### v1

# SkillDiffuser: Interpretable Hierarchical Planning via Skill Abstractions in Diffusion-Based Task Execution

Abstract
Diffusion models have demonstrated strong potential for robotic trajectory planning. However, generating coherent and long-horizon trajectories from high-level instructions remains challenging, especially for complex tasks requiring multiple sequential skills. We propose SkillDiffuser, an end-to-end hierarchical planning framework integrating interpretable skill learning with conditional diffusion planning to address this problem. At the higher level, the skill abstraction module learns discrete, human-understandable skill representations from visual observations and language instructions. These learned skill embeddings are then used to condition the diffusion model to generate customized latent trajectories aligned with the skills. This allows for generating diverse state trajectories that adhere to the learnable skills. By integrating skill learning with conditional trajectory generation, SkillDiffuser produces coherent behavior following abstract instructions across diverse tasks. Experiments on multi-task robotic manipulation benchmarks like Meta-World and LOReL demonstrate state-of-the-art performance and human-interpretable skill representations from SkillDiffuser. More visualization results and information could be found on our website.
1 Introduction
Figure 1: Comparison of SkillDiffuser and previous language conditioned diffusers. SkillDiffuser utilizes high-level abstraction to translate visual observations and language instructions into human understandable skills with language grounding. It then enables the low-level diffusion model condition on these skills, not only improving the execution performance of long-horizon, multi-step composite tasks but greatly enhancing the generalization and adaptability of the framework.
Recent research [17, 18, 6, 5] has demonstrated stronger generation abilities than previous generative models that help enhance reinforcement learning across various dimensions, including the generation of action trajectories [18, 1], representation of policies [48, 4], and synthesis of data [13, 22]. However, their ability to generate coherent trajectories for intricate tasks still poses challenges in terms of performance and generalizability, as these tasks require the completion of abstract instructions comprising numerous coordination-intensive sequential steps.
Previous approaches, [13, 1], such as Decision Diffuser, aim to tackle this challenge by decomposing complex tasks into simpler sub-skills, organized within a predefined skill library. These methods rely on conditioning the diffusion model with one-hot skill representations to generate trajectories for each of these sub-skills. Nevertheless, these techniques encounter difficulties in attempting to autonomously learn end-to-end from a wide range of datasets, which hinders their scalability and ability to achieve end-to-end learning. In addition, without explicitly learning reusable skills, models cannot capture intricate inter-step dependencies and constraints, yielding fragmented and illogical trajectories. Decomposing ambitious instructions into learnable sub-goals and skills can better enable models to follow logical step sequences, respect task structures, and transfer common procedures (e.g., reusable skill abstraction and adaptable skill-based diffusion) between different tasks. This skill-centric paradigm paves the way for diffusion models that can interpret and execute elaborate, abstract instructions necessitating numerous sequential steps.
Inspired by the above observations, we propose SkillDiffuser, a hierarchical planning framework unifying high-level skill learning with low-level conditional diffusion-based execution. As in Fig. 1, compared with previous language-conditioned diffusers, SkillDiffuser is able to interpret and execute complex instructions end-to-end with higher transferability. SkillDiffuser induces interpretable sub-latent goals by learning reusable skills tailored to the task instructions. The framework conditions a diffusion model on these learned skills to generate customized, coherent trajectories aligned with the overall objectives. By combining hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can produce consistent, skill-oriented behavior without the need for a predefined skill library and fully observed states. This end-to-end approach with learnable skills equips SkillDiffuser to follow abstract instructions across diverse tasks.
SkillDiffuser works as follows. 1) A skill predictor with vector quantization [43] is used for high-level skill learning to distill tasks into discrete and interpretable skills. Rather than forecasting skill durations, we adopt a fixed prediction horizon – predicting skills at regular intervals. This horizon-based discretization process seamlessly integrates visual and linguistic inputs into a cohesive skill set guiding the low-level diffusion model. 2) For skill-based trajectory generation, our diffusion model adopts a classifier-free conditioning approach, directly embedding skills into the model without additional classifiers. This allows for generating diverse state trajectories aligned with skill specifications while avoiding overfitting to a closed dataset. 3) To enable action inference from predicted states, we train an inverse dynamics network to decode motions between two consecutive frames generated. By separating state prediction from motion decoding, SkillDiffuser yields a fully adaptive framework for directing diverse embodiments via transferable state-space plans.
We evaluate the model’s performance in both skill learning and multi-task planning on the LOReL [28] Sawyer and Meta-World [49] datasets, with crucial experimental settings by considering robotic agents in real-world scenarios must operate with limited state information, chiefly visual observations. Furthermore, we present visualization results of SkillDiffuser, especially on skill learning to illustrate the model does have the ability to abstract high-level skills that are not only effective but also understandable to humans, bringing us closer to intelligent agents that acquire skills in an intuitive manner.
Our contributions are three-fold: 1) We propose an end-to-end hierarchical planning framework via skill learning for sub-goal abstraction; 2) We adopt a classifier-free diffusion model conditioned on learned skills to generate skill-oriented transferable state trajectories; 3) We demonstrate state-of-the-art performance on complex benchmarks and provide interpretable visualizations of human-understandable skill representations.
2 Related Works
2.1 Imitation Learning and Multi-task Learning
Imitation learning (IL) has evolved from foundational behavioral cloning to sophisticated multi-task learning frameworks. With traditional approaches relying on supervised learning from expert demonstrations [32, 36, 37], recent advancements have shifted towards learning the reward [15] or Q-function [11] from expert data, enhancing the ability to mimic complex behaviors. A new challenge lies in multi-task IL [39], where imitators are trained across varied tasks, aiming for generalization to new scenarios with task specifications ranging from vector states [27] to vision and language descriptions [12, 7, 46, 10].
Multi-task learning approaches often leverage shared representations to learn a spectrum of tasks simultaneously, enhancing the flexibility and efficiency of the learning process. The Meta-World benchmark [49] assesses multi-task and meta reinforcement learning, highlighting the need for algorithms capable of rapid adaptation. Building on this, the Prompting Decision Transformer [47] showcases few-shot policy generalization using task-specific prompts. Additionally, diffusion models have been explored as planners in multi-task settings [13], showing proficiency in generating diverse behaviors across tasks.
While these works represent significant strides in multi-task learning, the reliance on state-based inputs rather than pure visual inputs is a notable limitation, pointing to an important direction for future research.
2.2 Skill Discovery and Hierarchical Learning
Skill learning, fundamentally, referring to the process by which robots acquire new abilities or refine existing ones, is gaining increasing attention due to its pivotal role in enabling autonomous systems to adapt to new tasks, improve their performance over time, and interact more naturally with humans and complex environments. Traditionally, this domain was influenced by hand-crafted features and expert demonstrations [31].
With the development of deep learning, Eysenbach et al. [8] and A. Sharma et al. [38] investigated skill discovery in learning methods, achieving policies conditioned on learned latent variables and maintaining consistent skill codes throughout trajectories. In the domain of skill learning through language instructions, LISA [12] stands out by sampling multiple skills per trajectory, integrating language conditioning in a unique manner. Other studies have also explored hierarchical techniques in reinforcement learning [50, 26, 21]. These contributions enrich the field of skill discovery and hierarchical learning, offering diverse perspectives and methodologies.
2.3 Planning with Diffusion Model
Diffusion models [17] make great breakthroughs in image synthesis [17] recent years and has shown promising results in various generative applications. A seminal work that performs planning with diffusion directly is Diffuser [18], which laid the groundwork for using diffusion models in behavior synthesis. Then, a branch of this kind planning methods achieved state-of-the-art performance in a variety of decision tasks [30, 2, 6, 22, 4]. Among them, the work done by Chi et al. in [4] introduces the concept of learning the gradient of the action-distribution score function and iteratively optimizing it, demonstrating its significant potential in visuomotor policy learning. These works have further extended this direction and strengthened the versatility and generalization of diffusion-based planners.
Our approach is inspired by classifier-free diffusion guidance [16], which offers a significant advantage over traditional classifier-guided models [5]. By adopting the classifier-free approach, we can circumvent the challenges associated with training a reward model and $Q$-function, which are particularly cumbersome in many real-world planning scenarios of which the complexity is very high. Recent studies have also extended this direction, which use conditional diffusion models to generate customized trajectories. Decision Diffuser [1] is designed to generate trajectories for predefined skills. However, its limited one-hot skill encoding hinders its ability to autonomously learn from large datasets. This highlights the necessity for diffusion models with dynamic, learnable skill abstractions, facilitating robust multi-task learning without explicit state representations or pre-set skills. However, developing such an interpretable, end-to-end skill learning framework within diffusion models is still a challenge, and addressing these limitations is critical for enabling diffusion-based planners to effectively execute complex instructions through learned skill representations.
3 Preliminary
3.1 Planning with Diffusion over States
As introduced in previous works [18, 1], diffusion model is a promising tool to address the problem of planning in reinforcement learning which is cast as a Markov Decision Process (MDP) [33]. Within the MDP framework $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, the planning policy aims to identify an optimal action sequence $\boldsymbol{a}_{0:T}^{*}$ that maximizes the expected cumulative rewards over a finite time horizon $T$, governed by the state transition dynamics $\mathcal{T}$ and reward function $\mathcal{R}$. $\mathcal{S}$ is the state space and $\mathcal{A}$ is the action space.
By treating the state trajectory as sequence data $\boldsymbol{\tau}$, with sequence modeling, diffusion probabilistic models conceptualize planning as an iterative denoising process. The model progressively refines trajectories by reversing a forward diffusion process that is modeled as a Gaussian process, whereby noise is incrementally added to the data, denoted as $p_{\theta}(\boldsymbol{\tau}^{i-1}\mid\boldsymbol{\tau}^{i})$. Training involves minimizing the ELBO of the data’s negative log-likelihood, similar to variational Bayesian inference, with the optimization objective:
$\theta^{*}=\arg\min_{\theta}-\mathbb{E}_{\boldsymbol{\tau}^{0}}\left[\log p_{\theta}\left(\boldsymbol{\tau}^{0}\right)\right],\vspace{-5pt}$ (1)
where $p\left(\boldsymbol{\tau}^{N}\right)$ is a standard normal distribution and $\boldsymbol{\tau}^{0}$ denotes noiseless sequence data.
For practical implementation, a simplified surrogate loss function is proposed in [17], focusing on predicting the Gaussian mean of the reverse diffusion step:
$\mathcal{L}_{\text{denoise}}(\theta)=\mathbb{E}_{i,\boldsymbol{\tau}^{0}\sim q,\epsilon\sim\mathcal{N}}[||\epsilon-\epsilon_{\theta}(\boldsymbol{\tau}^{i},i)||^{2}].$ (2)
3.2 Classifier-free Diffusion Guidance
On the basis of unconditioned diffusion-based method, in the realm of offline reinforcement learning, a critical endeavor is to generate trajectories with the highest reward-to-go. With the flourishing development of conditional diffusion models [5], classifier-guided approaches embark on this by infusing specific trajectory information (encoded in $\boldsymbol{y}(\boldsymbol{\tau})$), such as the return $\mathcal{J}(\boldsymbol{\tau}^{0})$ or designated constraints, into the diffusion process:
$q(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i-1}),\;\;\;\;p_{\theta}(\boldsymbol{\tau}^{i-1}|\boldsymbol{\tau}^{i},\boldsymbol{y}(\boldsymbol{\tau})).\vspace{-3pt}$ (3)
With assumptions specified in [9], we have
$\boldsymbol{\tau}^{i-1}\sim\mathcal{N}(\mu_{\theta}+\alpha\Sigma\nabla_{\boldsymbol{\tau}}\log p\left(\boldsymbol{y}(\boldsymbol{\tau}^{i})\right),\Sigma),\vspace{-5pt}$ (4)
where $\alpha$ is a hyperparameter that adjusts the guidance strength, $\Sigma$ is the specified covariance of the noise and $\mu_{\theta}$ is the learned mean value of noise in unconditional diffusion.
However, the classifier-guided diffusion model requires an accurate estimation of guidance gradient $\nabla_{\boldsymbol{\tau}}\log p\left(\boldsymbol{y}(\boldsymbol{\tau})\right)$ based on the trajectory classifier $\boldsymbol{y}(\boldsymbol{\tau})$, which may not be feasible or may introduce a separate, complex dynamic programming procedure to estimate a $Q$-function in the training process.
Then, classifier-free guided models offer an alternative, where guidance gradient is eschewed, and instead trajectories with desirable characteristics, i.e. $\boldsymbol{y}(\boldsymbol{\tau})$, are emphasized without explicitly training a classifier. It augments the trajectory generation process with a guidance signal that amplifies the features of high-reward or optimal characteristics that have been implicitly present in the data. Mathematically, the noise to add during the reverse diffusion process is manipulated as:
$\hat{\epsilon}=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)+\omega(\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y},i)-\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)),$ (5)
where $\omega$ is the guidance scale, and $\varnothing$ represents the absence of guidance. Setting $\omega=0$ removes the classifier-free guidance towards an unconditional diffusion model, while a large value of $\omega$ strengthens the influence of the conditional information during trajectory generation.
Also, the loss function to minimize can be rewritten as,
[…]
where $\beta$ is a hyperparameter that controls the probability of dropping specific conditional information $\boldsymbol{y}(\boldsymbol{\tau})$. This dropout technique aids in enhancing sample diversity while maintaining context relevance.
After training the noise prediction model $\epsilon_{\theta}$ with the above $\mathcal{L}(\theta)$, the trajectory is sampled starting from Gaussian noise and progressively denoised with modified $\hat{\epsilon}$ through Eq. 5, employing re-parameterization technique.
In summary, through classifier-free guidance, we can modulate the trajectory sampling process to require the generated trajectories more aligned with the desired features represented by $\boldsymbol{y}(\boldsymbol{\tau})$. This process iteratively applies the conditioned noise model to refine the target trajectories that contain future states satisfying the constraints.
4 Methodology
Figure 2: Overall framework of SkillDiffuser. It’s a hierarchical planning model that leverages the cooperation of interpretable skill abstractions at the higher level and a skill conditioned diffusion model at the lower level for task execution in a multi-task learning environment. The high-level skill abstraction is achieved through a skill predictor and a vector quantization operation, generating sub-goals (skill set) that the diffusion model employs to determine the appropriate future states. Future states are converted to actions using an inverse dynamics model. This unique fusion enables a consistent underlying planner across different tasks, with the variation only in the inverse dynamics model.
4.1 Overview
Building upon the motivations discussed above, we present SkillDiffuser, an advanced methodological framework for robust multi-task learning across varied robots. This dynamic approach leverages the cooperation of skill learning at the higher level and a conditional diffusion model at the lower level. An overall framework is illustrated in Fig. 2. Notably, as we utilize language-based representations for abstracted skills, we make task execution based on diffusion interpretable and understandable to humans.
SkillDiffuser is designed to only require visual and language input for operation, without the need for explicit agent state information or predefined skill libraries. By abstracting complex tasks into skills (sub-goals), SkillDiffuser enables the diffusion model to work in concert with these sub-goals, ensuring adaptability and flexibility. This unique fusion enables the underlying planner to function consistently across different tasks, with only changes occurring in the inverse dynamic model, offering an unprecedented approach to task learning.
4.2 High Level Interpretable Skill Abstractions
In our SkillDiffuser framework, the high level interpretable skill abstraction module plays a crucial role in understanding and executing complex tasks. However, given the multi-task environments we consider, each task with only a single language instruction may be broken down into a sequence of sub-tasks or skills, which are not explicitly delineated within the instruction itself. Furthermore, suppose the offline training dataset is denoted as $\mathcal{D}$, it consists of trajectories derived from a sub-optimal policy for various tasks. A trajectory $\boldsymbol{\tau}=(l,\{\boldsymbol{i}_{t},\boldsymbol{a}_{t}\}_{t=1}^{T})$ includes language description $l$ and a sequence of image observations and actions $(\boldsymbol{i}_{t},\boldsymbol{a}_{t})$ over time steps $T$, with no reward labels attached. But the trajectories do not indicate the boundaries between sub-tasks as well, which thus requires the proposed methods capable of segmenting and interpreting the tasks into sub-goals in an unsupervised manner.
To address this challenge, we build a skill abstraction component upon a horizon-based skill predictor adapted from GPT-2 [34]. It is designed to parse and decompose tasks by fusing visual input and natural language instructions, alongside a Vector Quantization (VQ) sub-module that discretizes the learned skills into a skill set. The specifics of this component are illustrated below.
Firstly, as we utilize only images as robot state information, we transform the images into latent space features with a fixed image encoder such as R3M [29] before images are processed by the skill predictor. Mathematically, we denote the image encoder by $\Phi_{im}:\mathcal{I}\rightarrow\mathcal{R}^{I}$, where $\mathcal{I}$ represents the space of input images and $\mathcal{R}^{I}$, the resultant feature space, serves to condense the visual information into a form that is conducive for high-level semantics. Simultaneously, we use a language encoder to pre-process the natural language instructions, which is formalized as $\Phi_{lang}:L\rightarrow\mathcal{R}^{L}$, with $L$ being the space of language instructions and $\mathcal{R}^{L}$ the language feature space. The outputs of both encoders are then fed into the skill predictor, which operates to integrate these two modalities.
The skill predicting process is as follows: An image $\boldsymbol{i}_{t}\in\mathcal{I}$ at time step $t$ is encoded into a visual embedding $\boldsymbol{s}_{t}=\Phi_{im}(\boldsymbol{i}_{t})$. This embedding $\boldsymbol{s_{t}}$ is then input to the skill predictor, along with the language instruction $l\in L$, through the language encoder’s output $\boldsymbol{l}_{t}=\Phi_{lang}(l)$. The skill predictor, represented as $f:\mathcal{R}^{I}\times\mathcal{R}^{L}\rightarrow\mathcal{C}$, generates a skill code $\boldsymbol{\tilde{z}}$ by $\boldsymbol{\tilde{z}}=f(\boldsymbol{s}_{t},\boldsymbol{l}_{t})$ that encapsulates task’s requirements interpreted from the visual and language inputs.
After that, Vector Quantization [43] operation $\mathbf{q(\cdot)}$ is taken to transform these predicted skills into a low-dimensional discrete set $\mathcal{C}$. The discrete skill set contains $K$ skill embeddings $\left\{\boldsymbol{z}^{1},\boldsymbol{z}^{2},\ldots,\boldsymbol{z}^{K}\right\}$ which represent different potential skills. VQ operation is achieved by mapping the latent $\tilde{\boldsymbol{z}}$ to its closest entry of the skill set with skill vectors updated to be the moving average of the embeddings $\boldsymbol{z}$ closest to them, same as [43]. This process is as follows,
$\tilde{\boldsymbol{z}}=f\left(\Phi_{im}(\boldsymbol{i}_{t}),\Phi_{lang}(l)\right),\vspace{-2pt}$ (7)
$\boldsymbol{z}=\mathbf{q}(\tilde{\boldsymbol{z}})=\operatornamewithlimits{arg\,min}_{\boldsymbol{z}^{k}\in\mathcal{C}}\|\tilde{\boldsymbol{z}}-\boldsymbol{z}^{k}\|_{2}.\vspace{-3pt}$ (8)
VQ enforces each learnt skill $\boldsymbol{z}$ to lie in $\mathcal{C}$, which can be thought as learning $K$ prototypes or cluster centers for the language embeddings utilizing $k$-means [23] algorithm. This acts as a bottleneck, restricting the flow of language information and aiding in the learning of discrete skill codes. The back-propagation through the non-differentiable quantization operation is achieved by a straight-through gradient estimator, which simply copies the gradients and enables the model to be trained end-to-end.
We make each selected skill code $\boldsymbol{z}$ maintain across a designated number of time steps, called as horizon, subtly navigates the challenge of various sub-task durations by simply retaining the same skill code across multiple horizons. This strategy not only preserves the flexibility required for diverse task execution but also maintains the consistency of the model’s architecture.
Importantly, the discrete nature of the learned skill codes enhances the interpretability and controllability of the system’s behavior, as each skill code is associated with some human-understandable language phrases. An example is depicted in Fig. 4. Through this method, SkillDiffuser can not only learn to perform tasks based on language instructions but also achieve them in a human-interpretable way, allowing for a deeper understanding and control of the decision-making process of specific embodied agents.
Figure 3: SkillDiffuser’s low level skill-conditioned diffusion planning model. Notably, while the schematic here employs images to represent visual features for illustrative purposes, in actual implementation, both the input to and the sampling output of the diffusion model are state embeddings. The current observation is also the feature embedding of current visual observation.
4.3 Low Level Skill-conditioned Diffusion Planning
As discussed in Sec. 1, although existing methods (e.g. Decision Diffuser [1]) have proposed conditional diffusion models with skill constraints, they can only generate trajectories satisfying pre-defined skills, because the conditioning in such models is achieved by a classifier which determines whether a generated trajectory aligns with a specific skill or not. Therefore, they do not truly achieve a diffusion model that can condition on an arbitrary learned skill.
To overcome these constraints and enable diffusion models to plan over a learned continuous spectrum of skills, we propose an approach that leverages the classifier-free guided diffusion model with the skills embedded.
SkillDiffuser begins by employing a diffusion model to operate over the continuous skill space $\mathcal{Z}$ learned during high-level skill abstraction. We employ a U-Net to serve as the noise prediction model $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and guide it by in-context conditioning. More specifically, we firstly utilize a skill MLP (similar to point-wise feed-forward network [44]), denoted as $\Lambda$, to align skill features with denoising model. After that, we fuse the skill embeddings $\Lambda(\boldsymbol{z})$ into the state features throughout the residual blocks of U-Net. Details are depicted in Fig. 3. In this way, we make the diffusion model contextually modulate the influence of skill embeddings at each step. And following Eq. 5, we can synthesize future states attending to these given skills. This is a significant shift from previous static conditioning framework to a more dynamic and adaptive trajectory generation process.
Moreover, another important feature of our approach is its state-only diffusion characteristic, which eschews the direct generation of actions in diffusion model to facilitate more generalizable state predictions. And as a better alternative, we utilize another MLP, denoted as $\Psi(\cdot)$, to perform inverse dynamics after state generation to infer feasible actions that can achieve transitions between two continuous states. Additionally, we integrate observations from the current frame to provide more detailed information for motion prediction and achieve closed-loop control. This process is mathematically expressed as:
$\boldsymbol{a}_{t}=\Psi([\tilde{\boldsymbol{s}}_{t},\tilde{\boldsymbol{s}}_{t+1}],\boldsymbol{i}_{t})\quad\text{for}\quad t=0,\ldots,T-1,$ (9)
where $\tilde{\boldsymbol{s}}_{t}$ and $\tilde{\boldsymbol{s}}_{t+1}$ are consecutive observation embeddings within $\tilde{\boldsymbol{\tau}}$ generated by diffusion model, $\boldsymbol{i}_{t}$ is the current observation and $\boldsymbol{a}_{t}$ is the inferred action.
As the resulting action sequence $\{\boldsymbol{a}_{0},\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{T-1}\}$ derived from the generated states, it encapsulates the skills to execute the tasks, which allows for remarkable adaptability across multiple tasks. And when faced with a new task, we need only to replace the inverse dynamics model $\Psi(\cdot)$ specific to the new task’s kinematics, with the architecture and parameters of diffusion model unchanged. Such modularity ensures that the generative capabilities of SkillDiffuser are not task-specific but can be leveraged across a diverse range of tasks with varying dynamics, which mirrors the complex skill interplay and adaptability seen in human learning. A schematic diagram of low level skill-conditioned diffusion planning model is shown in Fig. 3.
4.4 Training the SkillDiffuser
We engineered a threefold loss function, with each meticulously designed and straightforward to meet the distinct needs of respective modules within SkillDiffuser.
For the task-specific inverse dynamics model, we utilize a behavior cloning loss [42] to train our inverse MLP which emulates expert actions from observed state transitions. This loss, denoted as $\mathcal{L}_{inv}$, is formulated as:
$\mathcal{L}_{inv}=\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\left[\left\|\boldsymbol{a}-\Psi(\tilde{\boldsymbol{s}},\tilde{\boldsymbol{s}}^{\prime},\boldsymbol{i})\right\|_{2}^{2}\right],$ (10)
where the notations are similar to Eq. 9.
Correspondingly, the other parts of the proposed model are all task-agnostic. For high-level skill abstraction module, we consider the vector quantization (VQ) loss to optimize the skill predictor. The VQ loss, $\mathcal{L}_{\mathrm{VQ}}$, encourages the skill predictor embeddings to adhere to the skill set vectors, enhancing the interpretability and consistency of the skill representations. Inspired by VQ-VAE [43], we have
$\mathcal{L}_{\mathrm{VQ}}=\mathbb{E}_{\boldsymbol{\tau}}\left[\|\mathbf{q}(\tilde{\boldsymbol{z}})-\tilde{\boldsymbol{z}}\|_{2}^{2}\right],$ (11)
where $\tilde{\boldsymbol{z}}$ follows Eq. 7.
[…]
Task Instruction Random LCBC [40] LCRL [19] Lang DT [12] LISA [12] SkillDiffuser
close drawer 52% 50% 58% 10% 100% 95 $\pm$ 3.2%
open drawer 14% 0% 8% 60% 20% 55 $\pm$ 13.3%
turn faucet left 24% 12% 13% 0% 0% 55 $\pm$ 9.3%
turn faucet right 15% 31% 0% 0% 30% 25 $\pm$ 4.4%
move black mug right 12% 73% 0% 20% 60% 18 $\pm$ 3.9%
move white mug down 5% 6% 0% 0% 30% 10 $\pm$ 1.7%
Average over tasks 20% 29% 13% 15% 40% 43 $\pm$ 1.1%
Table 1: Task-wise success rates on LOReL Sawyer Dataset. We show our success rates compared to random policy, language-conditioned imitation learning (LCBC), language-conditioned Q-learning (LCRL), a flat non-hierarchical Decision Transformer (Lang-DT), and LISA. The results on each dataset are calculated over 3 seeds. SkillDiffuser outperforms all other methods in terms of average performance over all tasks. Best methods and those within 5% of the best are highlighted in bold.
Rephrasal Type Lang DT LISA [12] SkillDiffuser
seen 15 40 43.65 ±$\pm 4.7$
unseen noun 13.33 33.33 36.01 ±$\pm 6.3$
unseen verb 28.33 30 36.70 ±$\pm 9.5$
unseen noun+verb 6.7 20 42.02 ±$\pm 3.8$
human provided 26.98 27.35 40.16 ±$\pm 2.1$
Average 18.07 30.14 39.71
Table 2: Rephrasal-wise success rates (in %) on LOReL Sawyer. Results of Lang DT, LISA and our SkillDiffuser are shown here. The standard error is calculated over 3 random seeds.
Additionally, the language encoder is achieved by leveraging a pre-trained GPT-2 [34] with parameters frozen to ensure stability in the language understanding component. And we employ diverse settings to serve as the visual encoder to ensure fair comparison in different datasets. We elaborate the details in corresponding parts of Sec. 5. More training details are shown in Appendix E and we also provide some pseudo-code of our algorithm in Appendix B.
5 Experiments
We first present a comprehensive evaluation on the LOReL Sawyer dataset, and then perform the ablation study compared on Meta-World benchmark to illustrate the efficiency of our method. Finally, we visualize the learned skills of our method both on LOReL and Meta-World MT10.
5.1 Datasets
LOReL Sawyer Dataset [28]. The Language-conditioned Offline Reward learning dataset, abbreviated as LOReL, is composed of pseudo-expert trajectories or play data gathered from the replay buffer of an arbitrary reinforcement learning policy, annotated with post-hoc crowd-sourced language directives. The inherent randomness within these trajectories presents a significant challenge for behavior cloning applications. The LOReL Sawyer dataset encompasses 50k trajectories, each with 20 steps, in a simulated Sawyer robot environment. We assess our approach using the same six tasks as the original paper [28]. This comes to a total of 77 instructions for all 6 tasks combined. More details about the dataset can be found in Appendix D.1.
Meta-World Dataset [49]. It is a comprehensive benchmark designed for evaluating multi-task and meta reinforcement learning algorithms. It introduces a comprehensive suite of 50 distinct robotic manipulation tasks, all located within a unified tabletop environment controlled by a simulated Sawyer arm. The Multi-Task 10 (MT10) within Meta-World is a subset comprising ten carefully selected tasks, balanced in terms of diversity and complexity. Details of the ten tasks can be found in Appendix D.2.
5.2 Evaluation Results on LOReL Sawyer Dataset
Settings. We first perform experiments on LOReL Sawyer dataset with only partially observed image data. The analysis is anchored on six different tasks with paraphrased instructions under five different situations (i.e. unseen verb, unseen none, human provided, etc. as introduced in Sec. 5.1). This deliberate choice facilitates a direct comparison of our method against established benchmarks.
Baselines. For our baselines, we employ random policy, language-conditioned imitation learning (LCBC) [40], language-conditioned Q-learning (LCRL) [19], Lang-DT (which is called Flat Baseline in [12]) and state-of-the-art skill-learning based method LISA [12]. We follow the settings of LOReL [28] paper for the first three algorithms, while follow LISA [12] for the last two.
The random policy serves as a baseline, executing actions without specific strategy or learning and providing a fundamental comparison point for the algorithm effectiveness. LCBC mimics offline dataset behaviors based on language instructions, aligning with previous works focusing on imitation learning to achieve language-conditioned behavior. In contrast, LCRL employs reinforcement learning, labeling each episode’s final state with language-instructed rewards. And Lang-DT serves as a non-hierarchical benchmark that utilizes a Causal Transformer [24] with language instructions. Moreover, LISA incorporates a dual-transformer structure, with one for skill prediction and the other for action planning, achieving state-of-the-art performance in language instructed multi-task learning.
Results. To ensure fair comparison, our method, SkillDiffuser, is designed to maintain a parameter count similar to baseline models. It employs the same visual and language encoder architecture as used in LISA [12] with meticulously matching of embedding dimensions and the number of heads across the layers of SkillDiffuser.
The results are present in Tables 1 and 2, showing task-wise and rephrasal-wise success rates for LOReL, averaged over 10 runs with a 20-step time horizon. SkillDiffuser, our approach, achieves the highest average performance in six different tasks, indicating its superior cross-task adaptability particularly when compared to similar approaches which yet are based on Decision Transformer [3], such as LISA [12]. Moreover, SkillDiffuser consistently excels in all rephrased types for LOReL test tasks, outperforming LISA by 9.6% on average. This demonstrates the model’s robustness in handling varied skill representations, marking a notable advancement in skill-conditioned diffusion model.
Method Lang Diffuser Skill Set Performance
Flat R3M [29] 13.3%
Lang Diffuser 16.7%
SkillDiffuser 23.3%
Table 3: Ablation study of language skill conditioning on Meta-World. Results of Flat R3M, Language-condition diffuser and our SkillDiffuser are shown here. All are averaged over 3 runs.
5.3 Ablation Study on Meta-World Dataset
Settings. We conduct experiments on Meta-World MT10 benchmark, a standard multi-task dataset which contains ten carefully selected tasks with sub-optimal expert data. We still do not use state data of the robotics but exploit visual observations only. We finely annotate the ten tasks with language instructions. Details are shown in Appendix D.2.
Baselines. We evaluate our methods against two baselines, both modifications of existing models. The first, Flat R3M, is adapted from the R3M paper [29]’s planner. As the original one utilizes the first four terms of robotic state information (i.e. proprio = 4 in R3M implementation), we reset “proprio” to zero, focusing exclusively on visual observations, and train the model on a multi-task dataset with 100 demonstrations per task (1000 total). To ensure fairness in comparison, we use R3M as our visual encoder in this setting. The second baseline is a variant of our SkillDiffuser, lacking the high-level skill abstraction module but retaining the low-level conditional diffusion-based planner, to assess the impact of skill abstraction. This version integrates a two-layer MLP to predict options from visual and language inputs, functioning as a language-conditioned diffuser, thus preserving the original setting’s core elements.
Ablation on Language Skill Conditioning. Table 3 indicates that the Meta-World MT10 tasks, reliant solely on visual inputs without state information, are challenging. The Flat R3M method [29], lacking language conditioning and skill sets, achieves moderate performance on tasks like drawer-close and reach just through behavior cloning. Traditional Lang Diffuser, with added language conditioning, outperforms Flat R3M. This suggests the value of language elements in enhancing performance. Unlike these diffusers that treat the entire task space as a single skill, our SkillDiffuser, by discretizing skills into a skill set, achieves a 6% higher performance than language-conditioned diffusers, demonstrating the effectiveness of skill acquisition.
5.4 Visualization Results of Skill Set
We show the visualization results of skill set on Meta-World Multi-Task 10 (MT10) here in Fig. 4 and result on LOReL dataset in Fig. 5 of Appendix C. The visual analysis of our SkillDiffuser’s learned skills on MT10 benchmark within Meta-World reveals that out of the 20-size skill set, our method learned eight skills (e.g. open door with revolving joint, pick and place goal etc.) notably distinguished by their unique word highlights. These bright spots across eight columns in the heatmap underscore the model’s ability to isolate and identify distinct skills from visual input, without an explicitly defined skill library. This indicates not only a significant interpretative advancement over traditional diffusion model-based planning but also the successful abstraction of high-level skill representations that are comprehensible to humans.
Figure 4: Visualization result of skill heat map on Meta-World Multi-Task 10 (MT10). We display the word frequency associated with a skill set of size 20 in Meta-World MT10, normalized by column. The data’s sparsity and distinct highlights indicate certain language tokens are uniquely linked to specific skills. There are eight skills learned by our method. (zoom in for best view)
6 Conclusion
This paper presents SkillDiffuser, an integrated framework that enables robots to execute tasks from natural language instructions by enabling interpretable skill learning and conditional diffusion planning. It uses vector quantization to learn discrete yet understandable skill representations directly from visual and linguistic demonstrations. These skills condition a diffusion model to generate state trajectories adhering to the learned skills. By combining hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can follow abstract instructions to complete manipulation tasks. Extensive Experiments on manipulation benchmarks demonstrate state-of-the-art performance, highlighting its effectiveness for long-horizon tasks and ability to automatically learn interpretable skills.
Supplementary Material
Appendix A Theoretical Foundation of Classifier-free Diffusion Model for Planning
A.1 Review of Classifier-guided Diffusion Model
Firstly, for a given trajectory $\boldsymbol{\tau}$, the standard reverse process of an unconditional diffusion probabilistic model is defined by $p_{\theta}(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i+1})$. This framework is then extended to incorporate conditioning on a specific label $\boldsymbol{y}$ (e.g., the reward), which is considered in the context of current-step denoised trajectory $\boldsymbol{\tau}^{i}$, which is represented as $p_{\phi}(\boldsymbol{y}|\boldsymbol{\tau}^{i})$. Consequently, the reverse diffusion process can be reformulated as $p_{\theta,\phi}(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i+1},\boldsymbol{y})$. This approach introduces additional parameters $\phi$ alongside the original diffusion model parameters $\theta$. The parameters $\phi$ can be viewed as a classifier, that encapsulates the probability of whether a noisy trajectory $\boldsymbol{\tau}^{i}$ satisfies the specific label $\boldsymbol{y}$, with a notation of $p_{\phi}(\boldsymbol{y}|\boldsymbol{\tau}^{i})$.
Under the constraints illustrated in [22, 5], we can derive the following theorem with lemma
$p_{\theta,\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i},\boldsymbol{\tau}^{i+1}\right)=p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right).$ (12)
Theorem A.1.
The conditional sampling probability of reverse diffusion process $p_{\theta,\phi}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1},\boldsymbol{y})$ is proportional to unconditional transition probability $p_{\theta}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1})$ multiplied by the classified probability $p_{\phi}(\boldsymbol{y}\mid\boldsymbol{\tau}^{i})$.
$p_{\theta,\phi}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1},\boldsymbol{y})=Zp_{\theta}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1})p_{\phi}(\boldsymbol{y}\mid\boldsymbol{\tau}^{i})$ (13)
Proof.
$\begin{split}p_{\theta,\phi}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1},&\ \boldsymbol{y})=\frac{p_{\theta,\phi}\left(\boldsymbol{\tau}^{i},\boldsymbol{\tau}^{i+1},\boldsymbol{y}\right)}{p_{\theta,\phi}\left(\boldsymbol{\tau}^{i+1},\boldsymbol{y}\right)}\\ &=\frac{p_{\theta,\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i},\boldsymbol{\tau}^{i+1}\right)p_{\theta}\left(\boldsymbol{\tau}^{i},\boldsymbol{\tau}^{i+1}\right)}{p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i+1}\right)p_{\theta}\left(\boldsymbol{\tau}^{i+1}\right)}\\ &=\frac{p_{\theta,\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i},\boldsymbol{\tau}^{i+1}\right)p_{\theta}\left(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1}\right)p_{\theta}\left(\boldsymbol{\tau}^{i+1}\right)}{p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i+1}\right)p_{\theta}\left(\boldsymbol{\tau}^{i+1}\right)}\\ &=\frac{p_{\theta,\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i},\boldsymbol{\tau}^{i+1}\right)p_{\theta}\left(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1}\right)}{p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i+1}\right)}\\ &=\frac{p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)p_{\theta}\left(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1}\right)}{p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i+1}\right)}.\end{split}$ (14)
The term $p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i+1}\right)$ is not directly correlated to $\boldsymbol{\tau}^{i}$ at the diffusion timestep $i$, thus can be viewed as a constant with notation $Z$. ∎
On this basis, using Taylor series expansion [14], we can sample trajectories by the modified Gaussian resampling.
Theorem A.2.
With a sufficiently large number of reverse diffusion steps, the sampling from reverse diffusion process $p_{\theta,\phi}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1},\boldsymbol{y})$ can be approximated by a modified Gaussian resampling. That is
$p_{\theta,\phi}(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i+1},\boldsymbol{y})\approx\mathcal{N}(\boldsymbol{\tau}^{i};\mu_{\theta}+\Sigma\nabla_{\boldsymbol{\tau}}\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right),\Sigma),$ (15)
where $\mu_{\theta}=\mu_{\theta}(\boldsymbol{\tau}^{i})$ and $\Sigma$ are the mean and variance of unconditional reverse diffusion process $p_{\theta}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1})$.
Proof.
With the above definition, we can rewrite the transfer probability of the unconditional denoising process as
$p_{\theta}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1})$ $=\mathcal{N}(\boldsymbol{\tau}^{i};\mu_{\theta},\Sigma)$ (16)
$\log p_{\theta}(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i+1})$ $=-\frac{1}{2}(\boldsymbol{\tau}^{i}-\mu_{\theta})^{T}\Sigma^{-1}(\boldsymbol{\tau}^{i}-\mu_{\theta})+C$ (17)
With a sufficiently large number of reverse diffusion steps, we apply Taylor expansion around $\boldsymbol{\tau}^{i}=\mu_{\theta}$ as
$\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)$ $=\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)|_{\boldsymbol{\tau}^{i}=\mu_{\theta}}$
$\hskip 10.0pt+\left.\left(\boldsymbol{\tau}^{i}-\mu_{\theta}\right)\nabla_{\boldsymbol{\tau}^{i}}\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)\right|_{\boldsymbol{\tau}^{i}=\mu_{\theta}}.$
Therefore, using Eq. 14, we derive
$\log p_{\theta,\phi}(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i+1},\boldsymbol{y})=\log p_{\theta}(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i+1})+\log p_{\phi}(\boldsymbol{y}|\boldsymbol{\tau}^{i})+C_{1}$
$\begin{split}RHS&=-\frac{1}{2}\left(\boldsymbol{\tau}^{i}-\mu_{\theta}\right)^{T}\Sigma^{-1}\left(\boldsymbol{\tau}^{i}-\mu_{\theta}\right)\\ &\hskip 10.0pt+\left(\boldsymbol{\tau}^{i}-\mu_{\theta}\right)\nabla\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)+C_{2}\end{split}$ (18)
$\begin{split}RHS&=-\frac{1}{2}\left(\boldsymbol{\tau}^{i}-\mu_{\theta}-\Sigma\nabla\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)\right)^{T}\times\Sigma^{-1}\\ &\hskip 10.0pt\times\left(\boldsymbol{\tau}^{i}-\mu_{\theta}-\Sigma\nabla\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right)\right)+C_{3},\end{split}$
which means,
$p_{\theta,\phi}(\boldsymbol{\tau}^{i}|\boldsymbol{\tau}^{i+1},\boldsymbol{y})\approx\mathcal{N}(\boldsymbol{\tau}^{i};\mu_{\theta}+\Sigma\nabla_{\boldsymbol{\tau}}\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right),\Sigma)$
A.2 Classifier-free Diffusion Model
While classifier guidance successfully achieves conditional guidance during trajectory generation, it is nonetheless reliant on gradients from a separate trained classifier which is hard to obtain in many cases. Classifier-free guidance [17] seeks to eliminate the classifier, which achieves the same effect as classifier guidance, but without such gradients.
First of all, we define the score function of the unconditional diffusion model as
$\epsilon_{\theta}(\boldsymbol{\tau}^{i})=-\Sigma\nabla_{\boldsymbol{\tau}}\log p_{\phi}\left(\boldsymbol{\tau}^{i}\right).$ (19)
Then, through Eq. 13 and 15, the score function of the classifier-guided diffusion model can be expressed as
$\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})=\epsilon_{\theta}(\boldsymbol{\tau}^{i})-\alpha\Sigma\nabla_{\boldsymbol{\tau}}\log p_{\phi}\left(\boldsymbol{y}\mid\boldsymbol{\tau}^{i}\right),$ (20)
where $\alpha$ is a scale hyper-parameter.
Theorem A.3.
Classifier-free guided diffusion model performs sampling with the linear combination of the conditional and unconditional score estimates as,
$\hat{\epsilon}_{\theta}=\hat{\epsilon}_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})=(1-\omega)\epsilon_{\theta}(\boldsymbol{\tau}^{i})+\omega\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y}),$ (21)
implicitly embedding guidance into the score function, with $\omega$ the scale hyper-parameter.
Proof.
Considering there is an implicit classifier denoted as $\tilde{p}_{\phi}(\boldsymbol{y}\mid\boldsymbol{\tau}^{i})$, with Bayes Rule [41], we can expand it as
$\tilde{p}_{\phi}(\boldsymbol{y}\mid\boldsymbol{\tau}^{i})\propto\tilde{p}_{\theta,\phi}(\boldsymbol{\tau}^{i},\boldsymbol{y})/p_{\theta}(\boldsymbol{\tau}^{i}).$
Then gradient of this implicit classifier would be
$\nabla_{\boldsymbol{\tau}}\log\tilde{p}_{\phi}(\boldsymbol{y}\mid\boldsymbol{\tau}^{i})=\nabla_{\boldsymbol{\tau}}\log\tilde{p}_{\theta,\phi}(\boldsymbol{\tau}^{i},\boldsymbol{y})-\nabla_{\boldsymbol{\tau}}\log p_{\theta}(\boldsymbol{\tau}^{i}).$ (22)
Substitute Eq. 19 in RHS, we get
$\alpha\Sigma\nabla_{\boldsymbol{\tau}}\log\tilde{p}_{\phi}(\boldsymbol{y}\mid\boldsymbol{\tau}^{i})$ $=\alpha\Sigma\nabla_{\boldsymbol{\tau}}\log\tilde{p}_{\theta,\phi}(\boldsymbol{\tau}^{i},\boldsymbol{y})$
$\hskip 60.0pt-\alpha\Sigma\nabla_{\boldsymbol{\tau}}\log p_{\theta}(\boldsymbol{\tau}^{i})$
$=-\alpha\hat{\epsilon}_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})+\alpha\epsilon_{\theta}(\boldsymbol{\tau}^{i})$
And then substitute Eq. 20 in LHS,
$\begin{split}\epsilon_{\theta}(\boldsymbol{\tau}^{i})-\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})&=-\alpha\hat{\epsilon}_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})+\alpha\epsilon_{\theta}(\boldsymbol{\tau}^{i})\\ \alpha\hat{\epsilon}_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})&=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})+(\alpha-1)\epsilon_{\theta}(\boldsymbol{\tau}^{i})\\ \hat{\epsilon}_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})=(1/&\alpha)\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})+(1-1/\alpha)\epsilon_{\theta}(\boldsymbol{\tau}^{i})\end{split}$ (23)
Let $\omega=1/\alpha$, we obtain,
$\hat{\epsilon}_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})=(1-\omega)\epsilon_{\theta}(\boldsymbol{\tau}^{i})+\omega\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y}),$
which is equal to Eq. 21. ∎
Therefore, in classifier-free diffusion guidance, we only need to train a single neural network to parameterize both conditional score estimator $\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y})$ and unconditional score estimator $\epsilon_{\theta}(\boldsymbol{\tau}^{i})$, where for the unconditional model we can set an empty set $\varnothing$ for the condition identifier $\boldsymbol{y}$ when predicting the score, i.e. $\epsilon_{\theta}(\boldsymbol{\tau}^{i})=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y}=\varnothing)$. Following the settings of [17], we jointly train the unconditional and conditional models simply by randomly setting $\boldsymbol{y}$ to the unconditional class identifier $\varnothing$ with probability $\beta$, which balances off the diversity and the relevance of the conditional label of generated samples.
Appendix B Pseudo-code of Training SkillDiffuser
As illustrated in Sec. 4.4, we provide the pseudocode for our SkillDiffuser’s training process in Algorithm 1, detailing its sequential stages and core mechanics. Additionally, Algorithm 2 describes the inference process, illustrating its steps of skill abstraction and trajectory generation.
Algorithm 1 Training process of SkillDiffuser
1:Dataset $\mathcal{D}$ of partially observed trajectories with paired language $\left\{\boldsymbol{\tau}_{\xi}=(l,\{\boldsymbol{i}_{t},\boldsymbol{a}_{t}\}_{t=0}^{T-1})\right\}_{\xi=1}^{N}$, size of the skill set $K$ and horizon $H$, pre-trained language and visual encoder $\Phi_{lang}$, $\Phi_{im}$
2:Initialize skill predictor $f$, conditional diffusion model $\mathcal{M}$, skill embedding model $\Lambda$ and inverse dynamics model $\Psi$
3:Vector Quantization op $\mathbf{q(\cdot)}$
4:while not converged do
5:       Sample $\tau=(l,\{\boldsymbol{i}_{t},\boldsymbol{a}_{t}\}_{t=0}^{T-1}$)
6:       Initialize partially observed states $S=\{\Phi(\boldsymbol{i}_{0})\}$
7:       for $k=0...\lfloor{\frac{T}{H}}\rfloor$ do $\triangleright$ Sample a skill every H steps
8:             $z\leftarrow\mathbf{q}(f(\Phi_{lang}(l),S))$
[…]
10:             for step $t=1...H$ do
11:                    $S\leftarrow S\cup\{\Phi(\boldsymbol{i}_{kH+t+1})\}$
12:                    $\tilde{\boldsymbol{a}}_{kH+t}\leftarrow\Psi([{\boldsymbol{s}}_{kH+t},{\boldsymbol{s}}_{kH+t+1}],\boldsymbol{i}_{kH+t})$triangleright Predict action using inverse dynamics model
13:                    $\mathcal{L}_{inv}=\mathbb{E}\left[\|\boldsymbol{a}_{kH+t}-\tilde{\boldsymbol{a}}_{kH+t}\|_{2}^{2}\right]$
14:                    Train $\Psi$ with objective $\mathcal{L}_{inv}$
15:             end for
[…]
17:       end for
18:end while
Algorithm 2 Inference process of SkillDiffuser
1:Initial partial observation $\boldsymbol{i}_{0}$ and the language instruction $l$, pre-trained language and visual encoder $\Phi_{lang}$, $\Phi_{im}$
2:Trained skill predictor $f$, conditional diffusion model $\mathcal{M}$, skill embedding model $\Lambda$ and inverse dynamics model $\Psi$
3:Initialize partially observed states $S=\{\Phi(\boldsymbol{i}_{0})\}$
4:for $k=0...\lfloor{\frac{T}{H}}\rfloor$ do $\triangleright$ Sample a skill every H steps
5:       $z\leftarrow\mathbf{q}(f(\Phi_{lang}(l),S))$
[…]
7:       for step $t=1...H$ do
8:             $\boldsymbol{a}_{kH+t}\leftarrow\Psi([\boldsymbol{s}_{kH+t},\boldsymbol{s}^{\prime}_{kH+t+1}],\boldsymbol{i}_{kH+t})$
9:             $\tilde{\boldsymbol{s}}_{kH+t+1}\leftarrow\text{Env.step}(\boldsymbol{a}_{kH+t})$triangleright Take action
10:             $S\leftarrow S\cup\{\tilde{\boldsymbol{s}}_{kH+t+1}\}$
11:       end for
12:end for
Figure 5: Visualization of skill heat map on LOReL. We display the word frequency associated with a skill set of size 20 in LOReL, normalized by column. The data’s sparsity and distinct highlights indicate certain language tokens are uniquely linked to specific skills. There are eleven skills learned by our method.
Appendix C More Visualizations
C.1 Visualization Results of Skill Set
Heat Map of Word Frequency in LOReL Dataset. As mentioned before, we show the visualization results of skill set on LOReL Sawyer Dataset in Fig. 5. The visualization results show that out of a 20-size skill-set, our SkillDiffuser learned 11 skills (e.g. shut close container drawer, pull drawer handle etc.) notably distinguished by their unique word highlights. The results demonstrate strong skill abstraction abilities. For example, the skill “shut close container drawer” abstracts different expressions like “shut drawer”, “shut container” into one skill semantic. In the heatmap, the presence of distinct bright spots across eleven columns strongly reaffirms the model’s capability to discern and pinpoint specific skills from visual inputs, in the absence of a pre-defined skill library. This observation is not just a testament to the model’s enhanced interpretative prowess over conventional diffusion-based planning approaches but also marks a remarkable stride in abstracting high-level skills into representations that are intuitively understandable by humans. Such evidence further validates the model’s proficiency in sophisticated skill identification and representation.
C.2 Word Cloud of Learned Skills
We further show the word cloud of 8 learned skills of LOReL Sawyer Dataset in Figure 6. From the results, we can find that the model has successfully mastered eight key skills, each closely linked to specific tasks. These skills demonstrate strong robustness to ambiguous language instructions. For instance, skill 4 effectively abstracts the skill of “open a drawer” from ambiguous expressions such as “open a container”, “pull a dresser”, “pull a drawer” and random combinations of these words. Similarly, skill 6 extracts the skill of “turn a faucet to the left”. This analysis indicates our method’s resilience to varied and poorly defined language inputs, confirming our SkillDiffuser can competently interpret and act upon a wide range of linguistic instructions, even those that are ambiguous or incomplete. These findings provide new perspectives and methodological guidance for future research in similar fields, especially in handling complex tasks with ambiguous language instructions. We also provide the word cloud of learned skills from Meta-World MT10 dataset in Fig. 7.
Figure 6: Word cloud of learned skills in LOReL Sawyer Dataset. We show eight of them here with the size corresponding to the word frequency in one skill.
Figure 7: Word cloud of learned skills in Meta-World MT10 Dataset. We show eight of them here with the size corresponding to the word frequency in one skill.
Task Description
Closing the Drawer
Involves the robot’s precise manipulation of a drawer to close it, testing spatial dynamics understanding and fine motor control.
Opening the Drawer Requires the robot to open a drawer, emphasizing its capability in tasks that necessitate pulling and spatial navigation.
Turning the Faucet Left
Assesses the robot’s precision in rotational movements for turning a faucet to the left, a nuanced everyday action.
Turning the Faucet Right Tests the robot’s adaptability in mirrored instructions, involving turning the faucet right, similar to the left turning task but in the opposite direction.
Pushing the Black Mug Right
Requires the robot to push a specific object (black mug) to the right, testing its skills in object recognition and directional movement.
Pushing the White Mug Down Involves pushing a different object (white mug) downward, further evaluating the robot’s ability to differentiate objects and execute varied motion commands.
Table 4: Overview of tasks in LOReL Sawyer Dataset.
Appendix D Dataset Descriptions
D.1 LOReL Sawyer Dataset
Figure 8: A sample instance of LOReL Sawyer Dataset. The start and goal images correspond to the instruction “close drawer”.
Language-conditioned Offline Reward Learning dataset, abbreviated as LOReL [28], contains trajectories originating from a reinforcement learning buffer which is generated by a random policy. The trajectories are sub-optimal and have language annotations through crowd-sourcing. Overall, the dataset encompasses approximately 50,000 language-annotated trajectories, each within a simulated environment featuring a Sawyer robot arm, with every demonstration extending over 20 discrete steps. A typical LOReL Sawyer environment is shown in Fig. 8. We assess our approach using the same set of instructions as those outlined in the original paper [28] which are described with their objectives in Tab. 4. These evaluation tasks are along with various rephrases of instructions which modify either the noun (“unseen noun”), the verb (“unseen verb”), both (“unseen noun+verb”), or entail a complete rewrite of the task (“human provided”), leading to a total of 77 distinct instructions for all six tasks. This structure of tasks and rephrases enables a comprehensive assessment of the robot’s ability to interpret and execute a wide range of language-based commands within the simulated environment.
D.2 Meta-World Dataset
The Meta-World dataset establishes a new benchmark in the field of multi-task and meta-reinforcement learning, offering 50 unique robotic manipulation tasks. These tasks range from simple to complex operations, providing researchers with a diverse testing ground. Each task is meticulously designed to ensure both challenge and common structural features that can be leveraged in multi-task and meta-learning algorithms. This design makes Meta-World an ideal choice for assessing the effectiveness and adaptability of algorithms in complex and variable task environments.
Particularly, the Multi-Task 10 (MT10) subset comprises 10 carefully selected tasks, where algorithms are trained and subsequently tested on the same set of tasks. As shown in Fig. 11, MT10 challenges algorithms’ learning and generalization capabilities in a multi-task environment, with the aim to evaluate the consistency and efficiency of algorithms in mastering multiple tasks, as well as their adaptability and robustness in the face of diverse tasks. As there is currently no widely-recognized instruction labeling of MT10, we provide our annotations here in Tab. 5.
We sample 100 trajectories for each task of MT10 and form the expert dataset of 1000 trajectories.
Task Identifier Language Instruction
window-close push and close a window
window-open push and open a window
door-open open a door with a revolving joint
peg-insert-side insert a peg sideways to the goal point
drawer-open open a drawer
pick-place pick a puck, and place the puck to the goal
reach reach the goal point
button-press-topdown press the button from the top
push push the puck to the goal point
drawer-close push and close a drawer
Table 5: Our annotated language instructions of Meta-World MT10 tasks.
Appendix E Implementation Details
E.1 Hyper-parameters
Generally, we follow the settings illustrated in [12] with details specified in the following Tab. 6.
Hyper-parameter LOReL Meta-World
Skill Predictor Transformer Layers 1 1
Skill Predictor Embedding Dim 128 128
Skill Predictor Transformer Heads 4 4
Skill Set Code Dim 16 16
Skill Set Size 20 20
Dropout 0.1 0.1
Batch Size 256 64
Skill Predictor Learning Rate 1e-6 1e-5
Conditional Diffuser Learning Rate 1e-3 5e-3
Inverse Dynamics Model Learning Rate 1e-3 5e-4
Diffuser Loss Weight 0.005 0.01
Horizon 8 8
VQ EMA Update 0.99 0.99
Skill Predictor and Diffuser Optimizer Adam Adam
Inverse Dynamics Model Optimizer Adam Adam
Table 6: Hyper-parameters of SkillDiffuser.
E.2 Architecture Details
  1. 1.
    We use 1 layer Transformer network for the skill predictor and follow the implementation of VQ-VAE [43] to achieve VQ operation.
  2. 2.
    The size of skill set is set to $20$ and the planning horizon is set to $8$ for all implementations.
  3. 3.
    A temporal U-Net [35] with 6 repeated residual blocks is employed to model the noise $\epsilon_{\theta}$ of the diffusion process. Each block is comprised of two temporal convolutions, each followed by group norm [45], and a final Mish non-linearity [25]. Timestep and skill embeddings are generated by two separate single fully-connected layer and added to the activation output after the first temporal convolution of each block.
Figure 9: Task-wise success rates (in %) on LOReL Sawyer Dataset.
Figure 10: Rephrasal-wise success rates (in %) on LOReL Sawyer Dataset.
E.3 Training Details
  1. 1.
    We train our model with one NVIDIA A100 Core Tensor GPU for about 45 hours in LOReL Sawyer dataset and about 24 hours in Meta-World MT10 dataset (1000 trajectories in total).
  2. 2.
    In both LOReL and Meta-World dataset, the skill predictor and diffusion model are trained with Adam optimizer [20] using a learning rate of $1\times 10^{-3}$ for the diffusion model, $1\times 10^{-6}$ for the LOReL skill predictor while $1\times 10^{-5}$ for Meta-World skill predictor. We only update parameters of Meta-World skill predictor every ten iterations. The inverse dynamics model is updated with Adam optimizer as well.
  3. 3.
    The batch size is set to $256$ for LOReL Sawyer dataset and $64$ for Meta-World MT10 dataset.
  4. 4.
    The training steps of the diffusion model are $5K$ for LOReL Sawyer dataset and $8K$ for Meta-World MT10 dataset. And the training epochs of the skill predictor are $500$ for both datasets.
  5. 5.
    The planning horizon $T$ of diffusion model is set to $100$ and the denoising steps are set to $200$ for all tasks.
Figure 11: Partially visual observations of all the 10 tasks in Meta-World MT10 Dataset.
Figure 12: Task-wise success rates (in %) on Meta-World MT10 Dataset.
Appendix F More Results
F.1 Task-wise Performance on LOReL Dataset
We further demonstrate the performance of our method and other baselines on LOReL Sawyer dataset in Fig 9 and 10. As can be seen from the figures, especially from Fig. 10, our method’s average performance on 5 rephrases is nearly 10 percentage points higher than the previous SOTA, which demonstrates its strong robustness against ambiguous language instructions.
F.2 Task-wise Performance on Meta-World Dataset
We also provide the task-wise success rates on Meta-World MT10 dataset in Fig. 12, achieved by Flat R3M [29], Language-conditioned Diffuser and SkillDiffuser. The average performance is shown separately in the right figure. From our experimental outcomes, it is clear to observe that our SkillDiffuser demonstrates commendable performance, particularly excelling in tasks involving mirrored instructions. SkillDiffuser exhibits an average performance enhancement of over 5% than previous language-conditioned Diffuser, which highlights the model’s advanced capability in understanding complex and ambiguous instructions compared to traditional methods. It showcases SkillDiffuser’s superior use of hierarchical architecture that employs interpretable skill learning for diffusion-based planners to better generate future trajectories.

### v3

@@
 Diffusion models have demonstrated strong potential for robotic trajectory planning.
-However, generating coherent and long-horizon trajectories from high-level instructions remains challenging, especially for complex tasks requiring multiple sequential skills.
+However, generating coherent trajectories from high-level instructions remains challenging, especially for long-range composition tasks requiring multiple sequential skills.
 We propose SkillDiffuser, an end-to-end hierarchical planning framework integrating interpretable skill learning with conditional diffusion planning to address this problem.
@@
 These learned skill embeddings are then used to condition the diffusion model to generate customized latent trajectories aligned with the skills.
-This allows for generating diverse state trajectories that adhere to the learnable skills.
+This allows generating diverse state trajectories that adhere to the learnable skills.
 By integrating skill learning with conditional trajectory generation, SkillDiffuser produces coherent behavior following abstract instructions across diverse tasks.
@@
 SkillDiffuser utilizes high-level abstraction to translate visual observations and language instructions into human understandable skills with language grounding.
-It then enables the low-level diffusion model condition on these skills, not only improving the execution performance of long-horizon, multi-step composite tasks but greatly enhancing the generalization and adaptability of the framework.
-Recent research [17, 18, 6, 5] has demonstrated stronger generation abilities than previous generative models that help enhance reinforcement learning across various dimensions, including the generation of action trajectories [18, 1], representation of policies [48, 4], and synthesis of data [13, 22].
-However, their ability to generate coherent trajectories for intricate tasks still poses challenges in terms of performance and generalizability, as these tasks require the completion of abstract instructions comprising numerous coordination-intensive sequential steps.
-Previous approaches, [13, 1], such as Decision Diffuser, aim to tackle this challenge by decomposing complex tasks into simpler sub-skills, organized within a predefined skill library.
+It then enables the low-level diffusion model condition on these skills, not only improving the execution performance of multi-step composition tasks but greatly enhancing the generalization and adaptability of the framework.
+Recent research [18, 19, 7, 6] has demonstrated diffusion models’ superior generative capabilities compared to previous models that help enhance reinforcement learning across various dimensions, including the generation of action trajectories [19, 2], policy representation [50, 5], and data synthesis [14, 23].
+However, their ability to generate coherent trajectories for intricate tasks still poses challenges in terms of performance and generalizability, as these tasks often require the fulfillment of abstract instructions that consist of numerous coordination-intensive sequential steps.
+Previous approaches [14, 2], such as Decision Diffuser, aim to tackle this challenge by decomposing complex tasks into simpler sub-skills, organized within a predefined skill library.
 These methods rely on conditioning the diffusion model with one-hot skill representations to generate trajectories for each of these sub-skills.
-Nevertheless, these techniques encounter difficulties in attempting to autonomously learn end-to-end from a wide range of datasets, which hinders their scalability and ability to achieve end-to-end learning.
+Nevertheless, these techniques encounter difficulties in attempting to autonomously learn end-to-end from a wide range of datasets, which hinders their scalability and ability to achieve end-to-end learning [1].
 In addition, without explicitly learning reusable skills, models cannot capture intricate inter-step dependencies and constraints, yielding fragmented and illogical trajectories.
-Decomposing ambitious instructions into learnable sub-goals and skills can better enable models to follow logical step sequences, respect task structures, and transfer common procedures (e.g., reusable skill abstraction and adaptable skill-based diffusion) between different tasks.
+Decomposing ambiguous instructions into learnable sub-goals and skills can better enable models to follow logical step sequences, respect task structures, and transfer common procedures (e.g.
+reusable skill abstraction and adaptable skill-based diffusion) between different tasks.
 This skill-centric paradigm paves the way for diffusion models that can interpret and execute elaborate, abstract instructions necessitating numerous sequential steps.
 Inspired by the above observations, we propose SkillDiffuser, a hierarchical planning framework unifying high-level skill learning with low-level conditional diffusion-based execution.
-As in Fig.
-1, compared with previous language-conditioned diffusers, SkillDiffuser is able to interpret and execute complex instructions end-to-end with higher transferability.
+As shown in Fig.
+1, compared with previous language-conditioned diffusion policies, SkillDiffuser is able to interpret and execute complex instructions end-to-end with higher transferability.
 SkillDiffuser induces interpretable sub-latent goals by learning reusable skills tailored to the task instructions.
 The framework conditions a diffusion model on these learned skills to generate customized, coherent trajectories aligned with the overall objectives.
-By combining hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can produce consistent, skill-oriented behavior without the need for a predefined skill library and fully observed states.
-This end-to-end approach with learnable skills equips SkillDiffuser to follow abstract instructions across diverse tasks.
+By integrating hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser achieves consistent, skill-oriented behavior without relying on a predefined skill library.
+Moreover, SkillDiffuser is designed to operate solely on visual observations, eliminating the need for robot proprioception (i.e.
+fully observed states).
+This end-to-end methodology, featuring learnable skills, enables SkillDiffuser to execute abstract instructions across a variety of tasks efficiently.
 SkillDiffuser works as follows.
@@
 This horizon-based discretization process seamlessly integrates visual and linguistic inputs into a cohesive skill set guiding the low-level diffusion model.
-2) For skill-based trajectory generation, our diffusion model adopts a classifier-free conditioning approach, directly embedding skills into the model without additional classifiers.
-This allows for generating diverse state trajectories aligned with skill specifications while avoiding overfitting to a closed dataset.
+2) For skill-based trajectory generation, we utilize a classifier-free diffusion model as policy, with skills directly embedded as guidance.
+This setting allows for generating multi-modal state trajectories aligned with skill specifications while avoiding overfitting to a closed dataset.
 3) To enable action inference from predicted states, we train an inverse dynamics network to decode motions between two consecutive frames generated.
@@
 We evaluate the model’s performance in both skill learning and multi-task planning on the LOReL [29] Sawyer and Meta-World [51] datasets, with crucial experimental settings by considering robotic agents in real-world scenarios must operate with limited state information, chiefly visual observations.
-Furthermore, we present visualization results of SkillDiffuser, especially on skill learning to illustrate the model does have the ability to abstract high-level skills that are not only effective but also understandable to humans, bringing us closer to intelligent agents that acquire skills in an intuitive manner.
+Furthermore, we present success rates on unseen compositional tasks, the reusability and visualizations of learned skills to illustrate the model does have the ability to abstract high-level skills that are not only effective but also understandable to humans, bringing us closer to intelligent agents that acquire skills in a direct manner.
 Our contributions are three-fold: 1) We propose an end-to-end hierarchical planning framework via skill learning for sub-goal abstraction; 2) We adopt a classifier-free diffusion model conditioned on learned skills to generate skill-oriented transferable state trajectories; 3) We demonstrate state-of-the-art performance on complex benchmarks and provide interpretable visualizations of human-understandable skill representations.
@@
 Building on this, the Prompting Decision Transformer [49] showcases few-shot policy generalization using task-specific prompts.
-Additionally, diffusion models have been explored as planners in multi-task settings [13], showing proficiency in generating diverse behaviors across tasks.
-While these works represent significant strides in multi-task learning, the reliance on state-based inputs rather than pure visual inputs is a notable limitation, pointing to an important direction for future research.
+And diffusion policy has also been explored in multi-task settings [14], which shows proficiency in generating diverse behaviors across tasks.
+However, unlike methods that leverage state inputs [14, 49] or access robot proprioception [30], SkillDiffuser uses raw visual inputs only.
 2.2 Skill Discovery and Hierarchical Learning
-Skill learning, fundamentally, referring to the process by which robots acquire new abilities or refine existing ones, is gaining increasing attention due to its pivotal role in enabling autonomous systems to adapt to new tasks, improve their performance over time, and interact more naturally with humans and complex environments.
+Skill learning, the process by which robots acquire new abilities or refine existing ones, is gaining increasing attention due to its pivotal role in enabling autonomous systems to adapt to new tasks, improve the performance over time, and interact naturally with humans and complex environments.
 Traditionally, this domain was influenced by hand-crafted features and expert demonstrations [32].
@@
 In the domain of skill learning through language instructions, LISA [13] stands out by sampling multiple skills per trajectory, integrating language conditioning in a unique manner.
-Other studies have also explored hierarchical techniques in reinforcement learning [50, 26, 21].
-These contributions enrich the field of skill discovery and hierarchical learning, offering diverse perspectives and methodologies.
+Our SkillDiffuser follows this way to extract sub-skills of each instruction at the higher-level.
+But differently, SkillDiffuser employs an adaptable diffusion policy at the lower-level to condition on different sub-skills to generate different actions, which formulates a creative hierarchical planning framework also advancing research on hierarchical techniques of reinforcement learning [52, 27, 22].
 2.3 Planning with Diffusion Model
@@
 These works have further extended this direction and strengthened the versatility and generalization of diffusion-based planners.
-Our approach is inspired by classifier-free diffusion guidance [16], which offers a significant advantage over traditional classifier-guided models [5].
+Our approach is inspired by classifier-free diffusion guidance [17], which offers a significant advantage over classifier-guided models [6].
 By adopting the classifier-free approach, we can circumvent the challenges associated with training a reward model and $Q$-function, which are particularly cumbersome in many real-world planning scenarios of which the complexity is very high.
 Recent studies have also extended this direction, which use conditional diffusion models to generate customized trajectories.
-Decision Diffuser [1] is designed to generate trajectories for predefined skills.
-However, its limited one-hot skill encoding hinders its ability to autonomously learn from large datasets.
-This highlights the necessity for diffusion models with dynamic, learnable skill abstractions, facilitating robust multi-task learning without explicit state representations or pre-set skills.
-However, developing such an interpretable, end-to-end skill learning framework within diffusion models is still a challenge, and addressing these limitations is critical for enabling diffusion-based planners to effectively execute complex instructions through learned skill representations.
+Decision Diffuser [2] is an example which is designed to generate trajectories for a predefined skill library.
+However, unlike our method, it can’t autonomously learn skill abstractions in an end-to-end fashion, which makes it difficult to scale to more tasks.
+This highlights the necessity for diffusion models with dynamic, learnable skill abstractions, facilitating complex instruction execution.
 3 Preliminary
@@
 where $\alpha$ is a hyperparameter that adjusts the guidance strength, $\Sigma$ is the specified covariance of the noise and $\mu_{\theta}$ is the learned mean value of noise in unconditional diffusion.
-However, the classifier-guided diffusion model requires an accurate estimation of guidance gradient $\nabla_{\boldsymbol{\tau}}\log p\left(\boldsymbol{y}(\boldsymbol{\tau})\right)$ based on the trajectory classifier $\boldsymbol{y}(\boldsymbol{\tau})$, which may not be feasible or may introduce a separate, complex dynamic programming procedure to estimate a $Q$-function in the training process.
-Then, classifier-free guided models offer an alternative, where guidance gradient is eschewed, and instead trajectories with desirable characteristics, i.e.
-$\boldsymbol{y}(\boldsymbol{\tau})$, are emphasized without explicitly training a classifier.
-It augments the trajectory generation process with a guidance signal that amplifies the features of high-reward or optimal characteristics that have been implicitly present in the data.
-Mathematically, the noise to add during the reverse diffusion process is manipulated as:
-$\hat{\epsilon}=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)+\omega(\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y},i)-\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)),$ (5)
+However, the classifier-guided diffusion model requires an accurate estimation of guidance gradient based on the trajectory classifier $\boldsymbol{y}(\boldsymbol{\tau})$, which may not be feasible and need to introduce a separate dynamic programming procedure to estimate a $Q$-function in the training process.
+Thus, classifier-free guidance offer an alternative, which augments the trajectory generation process with a guidance signal that amplifies the features of high-reward or optimal characteristics, i.e.
+$\boldsymbol{y}(\boldsymbol{\tau})$, that are implicitly present in the data.
+Mathematically, the noise to add during the reverse denoising process is:
+$\hat{\epsilon}=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)+\omega(\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y},i)-\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)),\vspace{-5pt}$ (5)
 where $\omega$ is the guidance scale, and $\varnothing$ represents the absence of guidance.
@@
 […]
-where $\beta$ is a hyperparameter that controls the probability of dropping specific conditional information $\boldsymbol{y}(\boldsymbol{\tau})$.
-This dropout technique aids in enhancing sample diversity while maintaining context relevance.
+where $\beta$ is a hyperparameter that controls the probability of dropping specific condition $\boldsymbol{y}(\boldsymbol{\tau})$, enhancing sample diversity while maintaining context relevance.
 After training the noise prediction model $\epsilon_{\theta}$ with the above $\mathcal{L}(\theta)$, the trajectory is sampled starting from Gaussian noise and progressively denoised with modified $\hat{\epsilon}$ through Eq.
@@
 4 Methodology
+4.1 Overview
+Building upon the motivations discussed above, we present SkillDiffuser, an advanced methodological framework for robust multi-task learning across various robots.
+This dynamic approach leverages the cooperation of skill learning at the higher level and a conditional diffusion model at the lower level.
+An overall framework is illustrated in Fig.
+2.
+Notably, we leverage language-grounded representations to abstract skills, thereby rendering the execution of tasks through our diffusion policy both interpretable and comprehensible to humans.
 Figure 2: Overall framework of SkillDiffuser.
@@
 This unique fusion enables a consistent underlying planner across different tasks, with the variation only in the inverse dynamics model.
-4.1 Overview
-Building upon the motivations discussed above, we present SkillDiffuser, an advanced methodological framework for robust multi-task learning across varied robots.
-This dynamic approach leverages the cooperation of skill learning at the higher level and a conditional diffusion model at the lower level.
-An overall framework is illustrated in Fig.
-2.
-Notably, as we utilize language-based representations for abstracted skills, we make task execution based on diffusion interpretable and understandable to humans.
-SkillDiffuser is designed to only require visual and language input for operation, without the need for explicit agent state information or predefined skill libraries.
-By abstracting complex tasks into skills (sub-goals), SkillDiffuser enables the diffusion model to work in concert with these sub-goals, ensuring adaptability and flexibility.
-This unique fusion enables the underlying planner to function consistently across different tasks, with only changes occurring in the inverse dynamic model, offering an unprecedented approach to task learning.
+Figure 3: SkillDiffuser’s low level skill-conditioned diffusion planning model.
+Notably, while the schematic here employs images to represent visual features for illustrative purposes, in actual implementation, both the input to and the sampling output of the diffusion model are state embeddings.
+The current observation is also the feature embedding of current visual observation.
 4.2 High Level Interpretable Skill Abstractions
@@
 The specifics of this component are illustrated below.
-Firstly, as we utilize only images as robot state information, we transform the images into latent space features with a fixed image encoder such as R3M [29] before images are processed by the skill predictor.
-Mathematically, we denote the image encoder by $\Phi_{im}:\mathcal{I}\rightarrow\mathcal{R}^{I}$, where $\mathcal{I}$ represents the space of input images and $\mathcal{R}^{I}$, the resultant feature space, serves to condense the visual information into a form that is conducive for high-level semantics.
+Firstly, as we utilize only images as robot state information, we transform the images into latent space features with a fixed image encoder (e.g.
+R3M [30]).
+For convenience, we denote the image encoder as $\Phi_{im}:\mathcal{I}\rightarrow\mathcal{R}^{I}$, where $\mathcal{I}$ represents the space of input images and $\mathcal{R}^{I}$, the resultant feature space, serves to condense the visual information into a form that is conducive for high-level semantics.
 Simultaneously, we use a language encoder to pre-process the natural language instructions, which is formalized as $\Phi_{lang}:L\rightarrow\mathcal{R}^{L}$, with $L$ being the space of language instructions and $\mathcal{R}^{L}$ the language feature space.
@@
 This process is as follows,
-$\tilde{\boldsymbol{z}}=f\left(\Phi_{im}(\boldsymbol{i}_{t}),\Phi_{lang}(l)\right),\vspace{-2pt}$ (7)
-$\boldsymbol{z}=\mathbf{q}(\tilde{\boldsymbol{z}})=\operatornamewithlimits{arg\,min}_{\boldsymbol{z}^{k}\in\mathcal{C}}\|\tilde{\boldsymbol{z}}-\boldsymbol{z}^{k}\|_{2}.\vspace{-3pt}$ (8)
-VQ enforces each learnt skill $\boldsymbol{z}$ to lie in $\mathcal{C}$, which can be thought as learning $K$ prototypes or cluster centers for the language embeddings utilizing $k$-means [23] algorithm.
+$\tilde{\boldsymbol{z}}=f\left(\Phi_{im}(\boldsymbol{i}_{t}),\Phi_{lang}(l)\right),\vspace{-12pt}$ (7)
+$\boldsymbol{z}=\mathbf{q}(\tilde{\boldsymbol{z}})=\operatornamewithlimits{arg\,min}_{\boldsymbol{z}^{k}\in\mathcal{C}}\|\tilde{\boldsymbol{z}}-\boldsymbol{z}^{k}\|_{2}.\vspace{-2pt}$ (8)
+VQ enforces each learnt skill $\boldsymbol{z}$ to lie in $\mathcal{C}$, which is equal to learning $K$ prototypes for the language embeddings utilizing $k$-means [24] algorithm.
 This acts as a bottleneck, restricting the flow of language information and aiding in the learning of discrete skill codes.
-The back-propagation through the non-differentiable quantization operation is achieved by a straight-through gradient estimator, which simply copies the gradients and enables the model to be trained end-to-end.
-We make each selected skill code $\boldsymbol{z}$ maintain across a designated number of time steps, called as horizon, subtly navigates the challenge of various sub-task durations by simply retaining the same skill code across multiple horizons.
-This strategy not only preserves the flexibility required for diverse task execution but also maintains the consistency of the model’s architecture.
+The back-propagation through the non-differentiable quantization is achieved by a straight-through gradient estimator, which simply copies the gradients and enables model to be trained end-to-end.
+In our approach, we apply a consistent skill code $\boldsymbol{z}$ across a defined number of time steps, termed the horizon.
+This consistent application across multiple horizons adeptly addresses the challenge of varying sub-task durations without altering the horizon itself.
+Consequently, this strategy not only preserves the flexibility required for diverse task execution but also maintains the model’s architectural stability by avoiding horizon-induced structural changes.
 Importantly, the discrete nature of the learned skill codes enhances the interpretability and controllability of the system’s behavior, as each skill code is associated with some human-understandable language phrases.
-An example is depicted in Fig.
-4.
+An example is depicted in Fig 6.
 Through this method, SkillDiffuser can not only learn to perform tasks based on language instructions but also achieve them in a human-interpretable way, allowing for a deeper understanding and control of the decision-making process of specific embodied agents.
-Figure 3: SkillDiffuser’s low level skill-conditioned diffusion planning model.
-Notably, while the schematic here employs images to represent visual features for illustrative purposes, in actual implementation, both the input to and the sampling output of the diffusion model are state embeddings.
-The current observation is also the feature embedding of current visual observation.
 4.3 Low Level Skill-conditioned Diffusion Planning
-As discussed in Sec.
-1, although existing methods (e.g.
-Decision Diffuser [1]) have proposed conditional diffusion models with skill constraints, they can only generate trajectories satisfying pre-defined skills, because the conditioning in such models is achieved by a classifier which determines whether a generated trajectory aligns with a specific skill or not.
-Therefore, they do not truly achieve a diffusion model that can condition on an arbitrary learned skill.
+As highlighted in Section 1, while existing approaches like Decision Diffuser [2] have introduced conditional diffusion models constrained by skills, their capability is limited to generating trajectories that meet only predefined skills.
+Consequently, these models fall short of realizing a diffusion model capable of conditioning on an arbitrary learned skill.
 To overcome these constraints and enable diffusion models to plan over a learned continuous spectrum of skills, we propose an approach that leverages the classifier-free guided diffusion model with the skills embedded.
@@
 This is a significant shift from previous static conditioning framework to a more dynamic and adaptive trajectory generation process.
-Moreover, another important feature of our approach is its state-only diffusion characteristic, which eschews the direct generation of actions in diffusion model to facilitate more generalizable state predictions.
-And as a better alternative, we utilize another MLP, denoted as $\Psi(\cdot)$, to perform inverse dynamics after state generation to infer feasible actions that can achieve transitions between two continuous states.
+Moreover, following previous works [2, 23, 7], we adopt a state-only diffusion model, which eschews the direct generation of actions in diffusion model.
+And instead, we utilize another MLP, denoted as $\Psi(\cdot)$, to perform inverse dynamics after state generation to infer feasible actions that can achieve transitions between two continuous states.
 Additionally, we integrate observations from the current frame to provide more detailed information for motion prediction and achieve closed-loop control.
-This process is mathematically expressed as:
-$\boldsymbol{a}_{t}=\Psi([\tilde{\boldsymbol{s}}_{t},\tilde{\boldsymbol{s}}_{t+1}],\boldsymbol{i}_{t})\quad\text{for}\quad t=0,\ldots,T-1,$ (9)
+Mathematically, it is:
+$\boldsymbol{a}_{t}=\Psi([\tilde{\boldsymbol{s}}_{t},\tilde{\boldsymbol{s}}_{t+1}],\boldsymbol{i}_{t})\quad\text{for}\quad t=0,\ldots,T-1,\vspace{-5pt}$ (9)
 where $\tilde{\boldsymbol{s}}_{t}$ and $\tilde{\boldsymbol{s}}_{t+1}$ are consecutive observation embeddings within $\tilde{\boldsymbol{\tau}}$ generated by diffusion model, $\boldsymbol{i}_{t}$ is the current observation and $\boldsymbol{a}_{t}$ is the inferred action.
 As the resulting action sequence $\{\boldsymbol{a}_{0},\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{T-1}\}$ derived from the generated states, it encapsulates the skills to execute the tasks, which allows for remarkable adaptability across multiple tasks.
-And when faced with a new task, we need only to replace the inverse dynamics model $\Psi(\cdot)$ specific to the new task’s kinematics, with the architecture and parameters of diffusion model unchanged.
-Such modularity ensures that the generative capabilities of SkillDiffuser are not task-specific but can be leveraged across a diverse range of tasks with varying dynamics, which mirrors the complex skill interplay and adaptability seen in human learning.
-A schematic diagram of low level skill-conditioned diffusion planning model is shown in Fig.
+And when faced with a new task, we are only required to change the inverse dynamics model $\Psi(\cdot)$ specific to the new task’s kinematics, with the architecture and parameters of diffusion model unchanged.
+Such modularity ensures the generative capabilities of SkillDiffuser are not task-specific but can be leveraged across a diverse range of tasks with varying dynamics.
+A schematic diagram of the low level module is shown in Fig.
 3.
 4.4 Training the SkillDiffuser
-We engineered a threefold loss function, with each meticulously designed and straightforward to meet the distinct needs of respective modules within SkillDiffuser.
-For the task-specific inverse dynamics model, we utilize a behavior cloning loss [42] to train our inverse MLP which emulates expert actions from observed state transitions.
+We engineered a threefold loss function for SkillDiffuser.
+Firstly, for the inverse dynamics model which is task-specific, we employ a behavior cloning loss [44] to train our inverse MLP emulating expert actions from observed state transitions.
 This loss, denoted as $\mathcal{L}_{inv}$, is formulated as:
-$\mathcal{L}_{inv}=\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\left[\left\|\boldsymbol{a}-\Psi(\tilde{\boldsymbol{s}},\tilde{\boldsymbol{s}}^{\prime},\boldsymbol{i})\right\|_{2}^{2}\right],$ (10)
+$\mathcal{L}_{inv}=\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\left[\left\|\boldsymbol{a}-\Psi(\tilde{\boldsymbol{s}},\tilde{\boldsymbol{s}}^{\prime},\boldsymbol{i})\right\|_{2}^{2}\right],\vspace{-5pt}$ (10)
 where the notations are similar to Eq. 9.
-Correspondingly, the other parts of the proposed model are all task-agnostic.
-For high-level skill abstraction module, we consider the vector quantization (VQ) loss to optimize the skill predictor.
-The VQ loss, $\mathcal{L}_{\mathrm{VQ}}$, encourages the skill predictor embeddings to adhere to the skill set vectors, enhancing the interpretability and consistency of the skill representations.
-Inspired by VQ-VAE [43], we have
-$\mathcal{L}_{\mathrm{VQ}}=\mathbb{E}_{\boldsymbol{\tau}}\left[\|\mathbf{q}(\tilde{\boldsymbol{z}})-\tilde{\boldsymbol{z}}\|_{2}^{2}\right],$ (11)
+Correspondingly, the other parts including both skill abstraction and low level execution are task-agnostic.
+For the high-level skill abstraction module, we apply a vector quantization (VQ) loss to refine the skill predictor.
+This VQ loss, $\mathcal{L}_{\mathrm{VQ}}$, ensures the embeddings produced by the skill predictor closely match the skill set vectors, thereby improving the interpretability and consistency of the skill representations.
+Inspired by VQ-VAE [45], we formulate it as:
+$\mathcal{L}_{\mathrm{VQ}}=\mathbb{E}_{\boldsymbol{\tau}}\left[\|\mathbf{q}(\tilde{\boldsymbol{z}})-\tilde{\boldsymbol{z}}\|_{2}^{2}\right],\vspace{-5pt}$ (11)
 where $\tilde{\boldsymbol{z}}$ follows Eq. 7.
@@
 […]
+Additionally, we utilize a pre-trained distilBERT [39] as our language encoder, adopting the configuration consistent with LOReL [29] and LISA [13], while freezing its parameters to guarantee stability in language understanding.
+And we employ diverse settings to serve as the visual encoder to ensure fair comparison in different datasets.
+We elaborate the details in corresponding parts of Sec.
+5.
+More training details are shown in Appendix G and we also provide some pseudo-code of our algorithm in Appendix B.
 Task Instruction Random LCBC [42] LCRL [20] Lang DT [13] LISA [13] SkillDiffuser
@@
 SkillDiffuser outperforms all other methods in terms of average performance over all tasks.
-Best methods and those within 5% of the best are highlighted in bold.
+Best methods and those within 6% of the best are highlighted in bold.
 Rephrasal Type Lang DT LISA [13] SkillDiffuser
@@
 Table 2: Rephrasal-wise success rates (in %) on LOReL Sawyer. Results of Lang DT, LISA and our SkillDiffuser are shown here. The standard error is calculated over 3 random seeds.
-Additionally, the language encoder is achieved by leveraging a pre-trained GPT-2 [34] with parameters frozen to ensure stability in the language understanding component.
-And we employ diverse settings to serve as the visual encoder to ensure fair comparison in different datasets.
-We elaborate the details in corresponding parts of Sec.
-5.
-More training details are shown in Appendix E and we also provide some pseudo-code of our algorithm in Appendix B.
 5 Experiments
@@
 5.1 Datasets
-LOReL Sawyer Dataset [28].
-The Language-conditioned Offline Reward learning dataset, abbreviated as LOReL, is composed of pseudo-expert trajectories or play data gathered from the replay buffer of an arbitrary reinforcement learning policy, annotated with post-hoc crowd-sourced language directives.
-The inherent randomness within these trajectories presents a significant challenge for behavior cloning applications.
+LOReL Sawyer Dataset [29] which is abbreviated from Language-conditioned Offline Reward learning, is composed of pseudo-expert trajectories or play data gathered from an arbitrary reinforcement learning policy, annotated with post-hoc crowd-sourced language directives.
 The LOReL Sawyer dataset encompasses 50k trajectories, each with 20 steps, in a simulated Sawyer robot environment.
-We assess our approach using the same six tasks as the original paper [28].
+We assess our approach using the same six tasks as the original paper [29] with paraphrased instructions under five different situations (i.e.
+seen, unseen verb, unseen none, unseen verb + noun and human provided).
 This comes to a total of 77 instructions for all 6 tasks combined.
@@
 5.2 Evaluation Results on LOReL Sawyer Dataset
-Settings.
-We first perform experiments on LOReL Sawyer dataset with only partially observed image data.
-The analysis is anchored on six different tasks with paraphrased instructions under five different situations (i.e.
-unseen verb, unseen none, human provided, etc.
-as introduced in Sec.
-5.1).
-This deliberate choice facilitates a direct comparison of our method against established benchmarks.
 Baselines.
-For our baselines, we employ random policy, language-conditioned imitation learning (LCBC) [40], language-conditioned Q-learning (LCRL) [19], Lang-DT (which is called Flat Baseline in [12]) and state-of-the-art skill-learning based method LISA [12].
-We follow the settings of LOReL [28] paper for the first three algorithms, while follow LISA [12] for the last two.
-The random policy serves as a baseline, executing actions without specific strategy or learning and providing a fundamental comparison point for the algorithm effectiveness.
-LCBC mimics offline dataset behaviors based on language instructions, aligning with previous works focusing on imitation learning to achieve language-conditioned behavior.
+We employ random policy, language conditioned imitation learning (LCBC) [42], language conditioned Q-learning (LCRL) [20], Lang-DT (also known as Flat Baseline in [13]) and state-of-the-art skill-learning method LISA [13] as our baselines.
+We follow the same settings as LOReL [29] for the first three algorithms, and follow LISA [13] for the last two.
+The random policy serves as a baseline.
+And LCBC mimics offline dataset behaviors based on instructions, aligning with previous works focusing on imitation learning to achieve language conditioned behavior.
 In contrast, LCRL employs reinforcement learning, labeling each episode’s final state with language-instructed rewards.
-And Lang-DT serves as a non-hierarchical benchmark that utilizes a Causal Transformer [24] with language instructions.
-Moreover, LISA incorporates a dual-transformer structure, with one for skill prediction and the other for action planning, achieving state-of-the-art performance in language instructed multi-task learning.
+Lang-DT plays as a non-hierarchical benchmark with a Causal Transformer [25], while LISA incorporates a dual-transformer structure, with one for skill prediction and the other for action planning.
+We do not compare with LOReL planner [29] as it uses MPC on a learned reward function relying on human annotations, while LISA and ours learn with trajectory data only.
 Results.
@@
 This demonstrates the model’s robustness in handling varied skill representations, marking a notable advancement in skill-conditioned diffusion model.
-Method Lang Diffuser Skill Set Performance
+Method Lang DT LOReL [29] LISA [13] SkillDiffuser
+Success Rate 13.33 ±$\pm 1.3$ 18.18 ±$\pm 1.8$ 20.89 ±$\pm 0.6$ 25.21 ±$\pm 2.7$
+Table 3: Performance on LOReL multi-step composition tasks.
+5.3 Performance on LOReL Compositional Tasks
+Settings.
+We conduct experiments following the same settings of unseen composition tasks of LISA [13] with 12 composition instructions in Tab.
+3.
+Detailed instructions are listed in Appendix D.2 with such an example that “open drawer and move black mug right”.
+We extend the max number of episode steps from customary 20 to 40, as LISA.
+Results.
+We observe SkillDiffuser achieves 2x the performance of non-hierarchical baseline (i.e.
+w/o skill abstraction) and also improves about 25% over LISA, highlighting its effectiveness.
+MPC-based LOReL planners are unable to perform as well in open scenarios like composition tasks.
+5.4 Ablation Study on Meta-World Dataset
+Settings. We conduct experiments on Meta-World MT10 benchmark with finely annotated instructions. We also use visual observations only. Details are in Appendix D.3.
+Baselines.
+We evaluate our method against three baselines, all modified from existing models.
+The first, Flat R3M, is adapted from R3M [30] paper’s planner.
+As the original one utilizes the first four terms of robot proprioception, we eliminate them and make the planner focus exclusively on visual observations.
+The second baseline is a variant of our SkillDiffuser, lacking the high-level skill abstraction module but retaining the low-level conditional diffusion-based planner, to assess the impact of skill abstraction.
+This version integrates a two-layer MLP to predict options from visual and language inputs, functioning as a language-conditioned diffusion planner.
+The last baseline is our re-implemented LISA [13] for Meta-World Dataset to validate the efficiency of diffusion model.
+To ensure fairness, we use R3M as the visual encoder for all of these methods on Meta-World.
+Method Lang Skill Set Diffuser Performance
 Flat R3M [30] 13.3%
+LISA [13] 13.8%
 Lang Diffuser 16.7%
 SkillDiffuser 23.3%
-Table 3: Ablation study of language skill conditioning on Meta-World. Results of Flat R3M, Language-condition diffuser and our SkillDiffuser are shown here. All are averaged over 3 runs.
-5.3 Ablation Study on Meta-World Dataset
-Settings.
-We conduct experiments on Meta-World MT10 benchmark, a standard multi-task dataset which contains ten carefully selected tasks with sub-optimal expert data.
-We still do not use state data of the robotics but exploit visual observations only.
-We finely annotate the ten tasks with language instructions.
-Details are shown in Appendix D.2.
-Baselines.
-We evaluate our methods against two baselines, both modifications of existing models.
-The first, Flat R3M, is adapted from the R3M paper [29]’s planner.
-As the original one utilizes the first four terms of robotic state information (i.e.
-proprio = 4 in R3M implementation), we reset “proprio” to zero, focusing exclusively on visual observations, and train the model on a multi-task dataset with 100 demonstrations per task (1000 total).
-To ensure fairness in comparison, we use R3M as our visual encoder in this setting.
-The second baseline is a variant of our SkillDiffuser, lacking the high-level skill abstraction module but retaining the low-level conditional diffusion-based planner, to assess the impact of skill abstraction.
-This version integrates a two-layer MLP to predict options from visual and language inputs, functioning as a language-conditioned diffuser, thus preserving the original setting’s core elements.
+Table 4: Ablation study of language skill conditioning on Meta-World. Results of Flat R3M, Language-condition diffuser, LISA and our SkillDiffuser are shown here. All are averaged over 3 runs.
 Ablation on Language Skill Conditioning.
-Table 3 indicates that the Meta-World MT10 tasks, reliant solely on visual inputs without state information, are challenging.
-The Flat R3M method [29], lacking language conditioning and skill sets, achieves moderate performance on tasks like drawer-close and reach just through behavior cloning.
-Traditional Lang Diffuser, with added language conditioning, outperforms Flat R3M.
-This suggests the value of language elements in enhancing performance.
-Unlike these diffusers that treat the entire task space as a single skill, our SkillDiffuser, by discretizing skills into a skill set, achieves a 6% higher performance than language-conditioned diffusers, demonstrating the effectiveness of skill acquisition.
-5.4 Visualization Results of Skill Set
-We show the visualization results of skill set on Meta-World Multi-Task 10 (MT10) here in Fig.
-4 and result on LOReL dataset in Fig.
-5 of Appendix C.
-The visual analysis of our SkillDiffuser’s learned skills on MT10 benchmark within Meta-World reveals that out of the 20-size skill set, our method learned eight skills (e.g.
-open door with revolving joint, pick and place goal etc.) notably distinguished by their unique word highlights.
-These bright spots across eight columns in the heatmap underscore the model’s ability to isolate and identify distinct skills from visual input, without an explicitly defined skill library.
-This indicates not only a significant interpretative advancement over traditional diffusion model-based planning but also the successful abstraction of high-level skill representations that are comprehensible to humans.
-Figure 4: Visualization result of skill heat map on Meta-World Multi-Task 10 (MT10).
-We display the word frequency associated with a skill set of size 20 in Meta-World MT10, normalized by column.
+Table 4 indicates our Meta-World MT10 task is quite different from and much more challenging than previous tasks that use states as observations [49] or take into robot proprioception [30].
+We only utilize single-frame visual input and instructions.
+The Flat R3M method, lacking language conditioning and skill sets can only succeed on tasks like drawer-close and reach through behavior cloning.
+Lang-conditioned Diffuser and modified LISA both outperform Flat R3M, suggesting the value of each corresponding module.
+Our SkillDiffuser, discretizing skills into a skill set, achieves a 6% higher performance than language-conditioned diffuser and a 9.5% higher than LISA, demonstrating the effectiveness of this combinational architecture.
+5.5 Ablation Study on Reusability of Learned Skills
+To evaluate the reusability of our learned skills, we calculate the average number of different skills used for a single instruction and the total number of instructions using each skill of LOReL Sawyer Dataset in Table 5.
+(With max episode step being 20, we experiment with skill horizon 10.) We observe each single instruction uses 1.55 sub-skills on average and each skill is called multiple times than the number of instructions (75 with 5 eval episodes), verifying the transferability of learnt skills.
+As suggested in Tab.10 of [13], except a very small horizon will hurt the performance, learning sub-skills to get refined semantics helps perform different actions at different stages.
+Besides, we also visualize resulting images from applying discrete skills in Appendix E.1 to further validate skills’ interpretability.
+5.6 Visualization Results of Learned Skill Set
+We show the visualization of skill set on LOReL compositional tasks here in Fig.
+4 and results on original LOReL dataset in Fig.
+5 and Meta-World dataset in Fig.
+6 in Appendix C.
+The visual analysis of our SkillDiffuser’s learned skills on LOReL compositional tasks reveals that out of the 20-size skill set, our method learned 11 skills (e.g.
+pull handle [skill 0], open counter [skill 14], etc.) notably distinguished by their unique word highlights.
+These bright spots across eleven columns (changing from only one column at initial which corresponds to default BC) in the heatmap underscore the model’s ability to identify and isolate distinct skills from visual inputs, without an explicitly defined skill library.
+This indicates not only a significant interpretative advancement over previous diffusion-based planning but a successful abstraction of high-level skill representations.
+# of learnt skills # of inst # of success Use 1 skill Use 2 skills Average
+17 375 144 64 80 1.55
+Freq of 17 skills 30, 8, 14, 10, 5, 19, 4, 7, 20, 2, 9, 20, 25, 6, 10, 3, 18
+Table 5: Average number of different skills used for a single instruction and total number of instructions used for each skill.
+Figure 4: Visualization of skill heat map on LOReL Sawyer compositional tasks.
+We display the word frequency associated with a skill set of size 20 in LOReL, normalized by column.
 The data’s sparsity and distinct highlights indicate certain language tokens are uniquely linked to specific skills.
-There are eight skills learned by our method.
+There are eleven skills learned by our method.
 (zoom in for best view)
 6 Conclusion
-This paper presents SkillDiffuser, an integrated framework that enables robots to execute tasks from natural language instructions by enabling interpretable skill learning and conditional diffusion planning.
-It uses vector quantization to learn discrete yet understandable skill representations directly from visual and linguistic demonstrations.
-These skills condition a diffusion model to generate state trajectories adhering to the learned skills.
-By combining hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can follow abstract instructions to complete manipulation tasks.
-Extensive Experiments on manipulation benchmarks demonstrate state-of-the-art performance, highlighting its effectiveness for long-horizon tasks and ability to automatically learn interpretable skills.
+This paper presents SkillDiffuser, an integrated framework that enables robots to perform tasks from natural language instructions by enabling interpretable skill learning and conditional diffusion planning.
+It employs vector quantization to learn discrete and comprehensible skill representations directly from visual and linguistic demonstrations.
+Subsequently, these skills condition a diffusion model to generate state trajectories adhering to the learned skills.
+Through integrating hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can comprehend and execute abstract instructions for various manipulation tasks.
+Extensive experiments on manipulation benchmarks demonstrate state-of-the-art performance, highlighting its effectiveness for multi-step composition tasks and ability to automatically learn interpretable skills.
+Acknowledgements
+This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622.
 Supplementary Material
@@
 There are eleven skills learned by our method.
+Figure 6: Visualization of skill heat map on Meta-World Multi-Task 10 (MT10). There are eight skills learned by our method. (zoom in for best view)
 Appendix C More Visualizations
-C.1 Visualization Results of Skill Set
-Heat Map of Word Frequency in LOReL Dataset.
+C.1 Visualization Results of Learned Skill Set
 As mentioned before, we show the visualization results of skill set on LOReL Sawyer Dataset in Fig.
-5.
-The visualization results show that out of a 20-size skill-set, our SkillDiffuser learned 11 skills (e.g.
-shut close container drawer, pull drawer handle etc.) notably distinguished by their unique word highlights.
+5 and Meta-World Multi-Task 10 (MT10) in Fig.
+6.
+The visualization results show that out of a 20-size skill-set, our SkillDiffuser learned 11 skills for LOReL (e.g.
+pull drawer handle [skill 0], shut close container drawer [skill 15], etc.) and 8 skills for Meta-World MT10 (e.g.
+open push window [skill 0], open door with revolving joint [skill 16], etc.).
 The results demonstrate strong skill abstraction abilities.
@@
 We also provide the word cloud of learned skills from Meta-World MT10 dataset in Fig.
-7.
+8.
 Figure 7: Word cloud of learned skills in LOReL Sawyer Dataset. We show eight of them here with the size corresponding to the word frequency in one skill.
 Figure 8: Word cloud of learned skills in Meta-World MT10 Dataset. We show eight of them here with the size corresponding to the word frequency in one skill.
-Task Description
-Closing the Drawer
-Involves the robot’s precise manipulation of a drawer to close it, testing spatial dynamics understanding and fine motor control.
-Opening the Drawer Requires the robot to open a drawer, emphasizing its capability in tasks that necessitate pulling and spatial navigation.
-Turning the Faucet Left
-Assesses the robot’s precision in rotational movements for turning a faucet to the left, a nuanced everyday action.
-Turning the Faucet Right Tests the robot’s adaptability in mirrored instructions, involving turning the faucet right, similar to the left turning task but in the opposite direction.
-Pushing the Black Mug Right
-Requires the robot to push a specific object (black mug) to the right, testing its skills in object recognition and directional movement.
-Pushing the White Mug Down Involves pushing a different object (white mug) downward, further evaluating the robot’s ability to differentiate objects and execute varied motion commands.
-Table 4: Overview of tasks in LOReL Sawyer Dataset.
 Appendix D Dataset Descriptions
@@
 Figure 9: A sample instance of LOReL Sawyer Dataset. The start and goal images correspond to the instruction “close drawer”.
+Task Description
+Closing the Drawer Involves the robot’s precise manipulation of a drawer to close it, testing spatial dynamics understanding and fine motor control.
+Opening the Drawer Requires the robot to open a drawer, emphasizing its capability in tasks that necessitate pulling and spatial navigation.
+Turning the Faucet Left Assesses the robot’s precision in rotational movements for turning a faucet to the left, a nuanced everyday action.
+Turning the Faucet Right Tests the robot’s adaptability in mirrored instructions, involving turning the faucet right, similar to the left turning task but in the opposite direction.
+Pushing the Black Mug Right Requires the robot to push a specific object (black mug) to the right, testing its skills in object recognition and directional movement.
+Pushing the White Mug Down Involves pushing a different object (white mug) downward, further evaluating the robot’s ability to differentiate objects and execute varied motion commands.
+Table 6: Overview of tasks in LOReL Sawyer Dataset.
 Language-conditioned Offline Reward Learning dataset, abbreviated as LOReL [29], contains trajectories originating from a reinforcement learning buffer which is generated by a random policy.
@@
 A typical LOReL Sawyer environment is shown in Fig.
-8.
+9.
 We assess our approach using the same set of instructions as those outlined in the original paper [29] which are described with their objectives in Tab.
-4.
+6.
 These evaluation tasks are along with various rephrases of instructions which modify either the noun (“unseen noun”), the verb (“unseen verb”), both (“unseen noun+verb”), or entail a complete rewrite of the task (“human provided”), leading to a total of 77 distinct instructions for all six tasks.
 This structure of tasks and rephrases enables a comprehensive assessment of the robot’s ability to interpret and execute a wide range of language-based commands within the simulated environment.
-D.2 Meta-World Dataset
-The Meta-World dataset establishes a new benchmark in the field of multi-task and meta-reinforcement learning, offering 50 unique robotic manipulation tasks.
-These tasks range from simple to complex operations, providing researchers with a diverse testing ground.
-Each task is meticulously designed to ensure both challenge and common structural features that can be leveraged in multi-task and meta-learning algorithms.
-This design makes Meta-World an ideal choice for assessing the effectiveness and adaptability of algorithms in complex and variable task environments.
-Particularly, the Multi-Task 10 (MT10) subset comprises 10 carefully selected tasks, where algorithms are trained and subsequently tested on the same set of tasks.
-As shown in Fig.
-11, MT10 challenges algorithms’ learning and generalization capabilities in a multi-task environment, with the aim to evaluate the consistency and efficiency of algorithms in mastering multiple tasks, as well as their adaptability and robustness in the face of diverse tasks.
-As there is currently no widely-recognized instruction labeling of MT10, we provide our annotations here in Tab.
-5.
-We sample 100 trajectories for each task of MT10 and form the expert dataset of 1000 trajectories.
+D.2 LOReL Composition Tasks
+We follow the same settings as LISA [13] to create 12 new composition tasks through combining original evaluation instructions as shown in Tab. 7.
+Instructions
+open drawer and move black mug right
+pull the handle and move black mug down
+move white mug right
+move black mug down
+close drawer and turn faucet right
+close drawer and turn faucet left
+turn faucet left and move white mug down
+turn faucet right and close drawer
+move white mug down and turn faucet left
+close the drawer, turn the faucet left and move black mug right
+open drawer and turn faucet counterclockwise
+slide the drawer closed and then shift white mug down
+Table 7: LOReL composition tasks
+Additionally, we also incorporate tasks such as “move white mug right” and “move black mug down” to explore the composition of skills related to colors (e.g., black and white) and directions (e.g., right and down).
+This aims to explore whether such skills can be combined to fulfill complex instructions.
 Task Identifier Language Instruction
@@
 drawer-close push and close a drawer
-Table 5: Our annotated language instructions of Meta-World MT10 tasks.
-Appendix E Implementation Details
-E.1 Hyper-parameters
+Table 8: Annotated instructions for Meta-World MT10 tasks.
+D.3 Meta-World Dataset
+The Meta-World dataset establishes a new benchmark in the field of multi-task and meta-reinforcement learning, offering 50 unique robotic manipulation tasks.
+These tasks range from simple to complex operations, providing researchers with a diverse testing ground.
+Each task is meticulously designed to ensure both challenge and common structural features that can be leveraged in multi-task and meta-learning algorithms.
+This design makes Meta-World an ideal choice for assessing the effectiveness and adaptability of algorithms in complex and variable task environments.
+Particularly, the Multi-Task 10 (MT10) subset comprises 10 carefully selected tasks, where algorithms are trained and subsequently tested on the same set of tasks.
+As shown in Fig.
+13, MT10 challenges algorithms’ learning and generalization capabilities in a multi-task environment, with the aim to evaluate the consistency and efficiency of algorithms in mastering multiple tasks, as well as their adaptability and robustness in the face of diverse tasks.
+As there is currently no widely-recognized instruction labeling of MT10, we provide our annotations here in Tab.
+8.
+We sample 100 trajectories for each task of MT10 and form the expert dataset of 1000 trajectories. We have released our dataset with image observations on https://skilldiffuser.github.io.
+Figure 10: Resulting images from applying skill 11 of Fig. 5. The black dashed line is a horizontal reference and please pay attention to the red oval region. (zoom in for best view)
+Appendix E More Ablations
+E.1 Ablation Study on Skill Interpretability
+Resulting Images from Applying Discrete Skills We visualize resulting images from applying skill 11 of Fig.
+5 which has grounding of “open, drawer, pull, dresser, container” (ranked from high frequency to low ones), consistent with its actual actions in Fig.
+10.
+We can clearly observe a behavior of pulling the drawer.
+And we would like to clarify not all skills have clear semantic or action correspondences, while some do.
+E.2 Ablation Study on Condition Guidance Weight
+Classifier-free guidance is widely used in generative model domain for its ability to act as temperature control when setting guidance weight above 1 during inference.
+In all of our experiments, we set guidance weight to 1.2 by default.
+But we also conduct ablation study on the condition guidance weight here in Tab.
+9.
+From the results, we find the guidance weight slightly greater than 1 helps the planner’s performance, while excessive weight hurts.
+Guidance Weight 1.0 1.2 1.8 3.0 5.0
+Success Rate on Seen Tasks 39.33% 46.67% 38.86% 39.03% 33.50%
+Table 9: Ablation on guidance weight. (5 episodes over 3 seeds.)
+Figure 11: Task-wise success rates (in %) on LOReL Sawyer Dataset.
+Figure 12: Rephrasal-wise success rates (in %) on LOReL Sawyer Dataset.
+Figure 13: Partially visual observations of all the 10 tasks in Meta-World MT10 Dataset.
+Figure 14: Task-wise success rates (in %) on Meta-World MT10 Dataset.
+Appendix F More Results
+F.1 Task-wise Performance on LOReL Dataset
+We further demonstrate the performance of our method and other baselines on LOReL Sawyer dataset in Fig 11 and 12.
+As can be seen from the figures, especially from Fig.
+12, our method’s average performance on 5 rephrases is nearly 10 percentage points higher than the previous SOTA, which demonstrates its strong robustness against ambiguous language instructions.
+F.2 Task-wise Performance on Meta-World
+We also provide the task-wise success rates on Meta-World MT10 dataset in Fig.
+14, achieved by Flat R3M [30], Language-conditioned Diffuser and SkillDiffuser.
+The average performance is shown separately in the right figure.
+From our experimental outcomes, it is clear to observe that our SkillDiffuser demonstrates commendable performance, particularly excelling in tasks involving mirrored instructions.
+SkillDiffuser exhibits an average performance enhancement of over 5% than previous language-conditioned Diffuser, which highlights the model’s advanced capability in understanding complex and ambiguous instructions compared to traditional methods.
+It showcases SkillDiffuser’s superior use of hierarchical architecture that employs interpretable skill learning for diffusion-based planners to better generate future trajectories.
+Appendix G Implementation Details
+G.1 Hyper-parameters
 Generally, we follow the settings illustrated in [13] with details specified in the following Tab. 10.
@@
 Conditional Diffuser Learning Rate 1e-3 5e-3
+Condition Guidance Weight 1.2 1.2
 Inverse Dynamics Model Learning Rate 1e-3 5e-4
@@
 Table 10: Hyper-parameters of SkillDiffuser.
-E.2 Architecture Details
+G.2 Architecture Details
   1. 1.
@@
 Timestep and skill embeddings are generated by two separate single fully-connected layer and added to the activation output after the first temporal convolution of each block.
-Figure 9: Task-wise success rates (in %) on LOReL Sawyer Dataset.
-Figure 10: Rephrasal-wise success rates (in %) on LOReL Sawyer Dataset.
-E.3 Training Details
+G.3 Training Details
   1. 1.
@@
     The planning horizon $T$ of diffusion model is set to $100$ and the denoising steps are set to $200$ for all tasks.
-Figure 11: Partially visual observations of all the 10 tasks in Meta-World MT10 Dataset.
-Figure 12: Task-wise success rates (in %) on Meta-World MT10 Dataset.
-Appendix F More Results
-F.1 Task-wise Performance on LOReL Dataset
-We further demonstrate the performance of our method and other baselines on LOReL Sawyer dataset in Fig 9 and 10.
-As can be seen from the figures, especially from Fig.
-10, our method’s average performance on 5 rephrases is nearly 10 percentage points higher than the previous SOTA, which demonstrates its strong robustness against ambiguous language instructions.
-F.2 Task-wise Performance on Meta-World Dataset
-We also provide the task-wise success rates on Meta-World MT10 dataset in Fig.
-12, achieved by Flat R3M [29], Language-conditioned Diffuser and SkillDiffuser.
-The average performance is shown separately in the right figure.
-From our experimental outcomes, it is clear to observe that our SkillDiffuser demonstrates commendable performance, particularly excelling in tasks involving mirrored instructions.
-SkillDiffuser exhibits an average performance enhancement of over 5% than previous language-conditioned Diffuser, which highlights the model’s advanced capability in understanding complex and ambiguous instructions compared to traditional methods.
-It showcases SkillDiffuser’s superior use of hierarchical architecture that employs interpretable skill learning for diffusion-based planners to better generate future trajectories.